### 14 Mar 2026 v1

# Fully distributed consensus control for stochastic multi-agent systems under undirected and directed topologies

Abstract
This work aims to address the design of fully distributed control protocols for stochastic consensus, and, for the first time, establishes the existence and uniqueness of solutions for the path-dependent and highly nonlinear closed-loop systems under both undirected and directed topologies, bridging a critical gap in the literature. For the case of directed graphs, a unified fully distributed control protocol is designed for the first time to guarantee mean square and almost sure consensus of stochastic multi-agent systems under directed graphs. Moreover, an enhanced fully distributed protocol with additional tunable parameters designed for undirected graphs is proposed, which guarantees stochastic consensus while achieving superior convergence speed. Additionally, our work provides explicit exponential estimates for the corresponding convergence rates of stochastic consensus, elucidating the relationship between the exponential convergence rate and the system parameters. Simulations validate the theoretical results.
I Introduction
In recent years, researchers have discovered that individual agents with limited communication and perception capabilities can effectively collaborate with each other in teams to accomplish more complex and diverse tasks that exceed their individual capabilities, and to some extent, save resources and costs [1, 2]. Consequently, the distributed collaborative control of multi-agent systems (MASs) has found widespread applications in various fields, including geological and marine exploration, agricultural automation, as well as military and aerospace [3, 4, 5, 6, 7]. Among these applications, the consensus problem [8] stands out as a typical and fundamental issue in distributed collaborative control.
Nevertheless, the design of control protocols discussed in the above works relies on Laplacian matrix information, which is a global information about the structure of communication topologies. This reliance imposes a significant limitation by hindering fully distributed control protocols and consequently undermining some advantages inherent to distributed control approaches. To overcome this constraint, Li et al. proposed for the first time a fully distributed protocol for consensus and tracking control of MASs, marking a significant shift toward local-information-based strategies [9]. Since then, researchers have made substantial progress in more complex situations through numerous exemplary contributions in the area of fully distributed control of deterministic MASs [10, 11].
A key development in this field is the widespread adoption of adaptive gain design, which enables dynamic control adjustment using only local information. Recent studies demonstrate this approach across diverse network topologies and system models. For instance, in [12], the consensus problem of second-order MASs with heterogeneous unknown time-varying inertias and control gains was studied under directed graphs through adaptive $\sigma$-modification schemes. [13] solved the problem of fully distributed consensus control for MASs with general linear dynamics under undirected graphs via adaptive dynamic event-triggered communication strategies. Further extending to signed networks, [14] proposes a time-varying parametric Lyapunov-based adaptive protocol for prescribed-time bipartite synchronization in cooperative-antagonistic networks. Studies such as [15] and [16] respectively ensure practical fixed-time consensus for single-integrator systems under undirected graphs and resolve leaderless/leader-following consensus for disturbed second-order systems under directed graphs, both employing distributed adaptive mechanisms. Moreover, [17] advances further by transcending topological and model constraints, proposing a topology-agnostic, data-driven event-triggered adaptive learning algorithm that achieves cooperative control under model uncertainty and communication constraints.
Collectively, these works exemplify the core idea of utilizing localized adaptive mechanisms to effectively operate across diverse network structures, ranging from undirected and directed graphs to signed digraphs, and progressing toward topology-agnostic scenarios, thereby illustrating an evolutionary trajectory from reliance on specific structural information to generalization independent of network topology. However, It should be noted that the aforementioned protocols primarily focus on deterministic MASs, leaving the challenge of stochastic disturbances largely unaddressed in current fully distributed control frameworks.
In modern engineering systems, interference is common and inevitable, and is caused by many factors such as complex communication environments, abrupt changes in working conditions, aging or damage of equipment, friction, and so on. Ignoring interference often results in an inability to meet high-precision control requirements in most practical applications. Therefore, considering the stochastic noise interference in the dynamics or communication topology of MASs in complex environments, researchers typically employ $\mathrm{It}\hat{\mathrm{o}}$ stochastic differential equations to model the dynamic behavior of each agent[18, 19, 20]. In recent years, there has been a wealth of outstanding research addressing stochastic MASs, and many studies in this area can be found in [21, 22, 23].
However, as far as we know, research on stochastic fully distributed consensus appears to be quite limited in the existing literature. In [24], the synchronization problem of a stochastic coupled nonlinear dynamic complex network under adaptive proportional-integral control was investigated. In [25], a fully distributed tracking control protocol was proposed for the stochastic nonlinear stochastic MASs with Markov switching topology by designing intermittent adaptive gains. Additionally, in [26], the fully distributed bipartite time-varying formation tracking control problem for heterogeneous linear MASs with stochastic disturbances was investigated under a signed Markovian switching topology. However, the aforementioned literature primarily focuses on synchronization or tracking control problems in the mean square(m.s.) sense and the interaction topologies among the nodes or followers are assumed to be undirected. The conventional analytical techniques for undirected graphs are proven inadequate for the case of directed graphs due to their reliance on symmetric properties–a fundamental feature absent in directed networks. This methodological incompatibility becomes particularly pronounced in stochastic cases, where two critical challenges arise: i) the existence of solutions for stochastic systems under fully distributed protocols requires rigorous verification, and ii) the almost sure(a.s.) and m.s. consensus in stochastic environment need theoretical examination. These inherent complexity of conducting consensus analysis and protocol synthesis for directed graphs under stochastic conditions leaves this a challenging problem. At present, to the best of our knowledge, for stochastic MASs, no existing work has addressed the design of fully distributed control protocols for directed graphs while ensuring stochastic a.s. consensus, which is one of the motivations for our work.
Notably, the implementation of fully distributed control protocols leads to path-dependent and highly nonlinear closed-loop dynamics [27], since the feedback involves integral functionals of the state history, which poses significant challenges for consensus analysis in stochastic MASs. For this case, it is necessary to prove the existence and uniqueness of the solution of closed-loop stochastic system in both undirected and directed graphs. However, although there have been some studies on fully distributed control of stochastic MASs under undirected graphs, no prior work has explored this fundamental issue, which is another motivation for this work.
Based on the preceding analysis, the contributions of this work are summarized as follows:
i) A unified fully distributed protocol is proposed to solve the m.s. and a.s. consensus problems for stochastic MASs under directed topology. To the best of our knowledge, this represents the first systematic study addressing fully distributed control for stochastic MASs under directed graphs, filling a notable gap in existing literature.
ii) For stochastic MASs under undirected graphs, an improved fully distributed protocol with rigorous convergence guarantees is developed. In contrast to prior works on fully distributed stochastic consensus, our analysis provides explicit exponential convergence rate estimates for both m.s. and a.s. consensus, establishing clear relationships between convergence performance and system parameters.
iii) By employing advanced stochastic analysis tools including the functional $\mathrm{It\hat{o}}$ formula and stopping-time techniques, we rigorously establish the existence and uniqueness of solutions for the closed-loop stochastic MASs (path-dependent stochastic systems), which feature inherent high nonlinearity under our proposed fully distributed protocols.
This paper is organized as follows. The problem formulation is provided in Section II. Section III presents the existence and uniqueness of the solution to the path-dependent stochastic systems. The main results on the fully distributed protocol design for stochastic MASs are presented in Section IV. In Section V, two simulations are provided to validate the theoretical findings. Section VI summarizes the paper.
Notations: $P^{T}$ represents the transposition of the $P$ matrix. $\mathbb{R}^{n\times m}$ is the set of all $n\times m$ real matrices. $P>0$ for $P\in\mathbb{R}^{n\times n}$ means that $P$ is a positive definite matrix. $\lambda_{\min}(P)$ and $\lambda_{\max}(P)$ represent the minimum and maximum eigenvalues of $P$. $I_{n}$ denotes the $n$-dimensional identity matrix. Let $(\Omega,{\mathcal{F}},\mathbb{P})$ denote a complete probability space with a filtration $\{{\mathcal{F}}_{t}\}_{t\geq 0}$ satisfying the usual conditions. For $\in R, pqp\wedge q$ and $p\vee q$ represent $\min\left\{p,q\right\}$ and $\max\left\{p,q\right\}$. $\sigma_{\max}(A)$ denotes the maximum singular value of a matrix $A\in\mathbb{R}^{m\times n}$. $\mathbf{1}_{N}$ denotes the $N$-dimensional column vector with all ones.
II Problem Formulation
[…]
Consider the following stochastic MASs with $N$ nodes where the dynamic of each agent is modeled by
$\begin{split}dx_{i}(t)=[Ax_{i}(t)+Bu_{i}(t)]dt+Cx_{i}(t)dw(t),\end{split}$ (1)
where $x_{i}(t)\in\mathbb{R}^{n}$ and $u_{i}(t)\in\mathbb{R}^{m}$ are the state and control input of the $i$th agent, respectively. i=1,2,,Ni=1,2,...,N, $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times m}$ and $C\in\mathbb{R}^{n\times n}$ are constant matrices, and $w(t)$ is a standard Brownian motion defined on the complete probability space $(\Omega,{\mathcal{F}},\mathbb{P})$. Let $x(t)=[x_{1}^{T}(t),...,x_{N}^{T}(t)]^{T}$ and $u(t)=[u_{1}^{T}(t),...,u_{N}^{T}(t)]^{T}$.
The unified fully distributed protocol is given as
$\begin{split}&u_{i}(t)=c_{i}(t)\Sigma_{i}(t)\mathcal{K}\xi_{i}(t),\\ &\dot{c}_{i}(t)=e^{\gamma t}\xi_{i}^{T}(t)\Gamma\xi_{i}(t),\end{split}$ (2)
where $\xi_{i}(t)=\sum_{j\in N_{i}}a_{ij}(x_{i}(t)-x_{j}(t))$, $c_{i}(t)\in\mathbb{R}$ denotes the time-varying adaptive gain, $c_{i}(0)>0$, and $\Sigma_{i}(t)\in\mathbb{R}$ represents the auxiliary time-varying gain, which is designed as
$\Sigma_{i}(t)=k_{1}\Big(k_{2}+\frac{\sigma_{i}(t)}{c_{i}(t)}\Big)^{\mu},$
with $\sigma_{i}(t)=\xi_{i}^{T}(t)P\xi_{i}^{T}(t)$, $\mu\geq 1,k_{1},k_{2}>0$. The feedback gain matrices $\mathcal{\mathcal{K}}\in\mathbb{R}^{m\times n}$, $\Gamma\in\mathbb{R}^{n\times n}$, $P\in\mathbb{R}^{n\times n}$ and constant $\gamma$ will be designed later in Section IV.
The goal is to design the control input $u(t)$ such that the consensus of the stochastic MASs (1) can be solved in m.s. or a.s. sense. The corresponding definitions are as follows.
Definition II.1.
The m.s. (or a.s.) consensus of the stochastic MASs (1) can be solved by the protocol $u(t)$ if for any initial value $x(0)\in\mathbb{R}^{nN}$ and all distinct $i,j\in\mathcal{V}$, $\lim_{t\to\infty}\mathbb{E}\|x_{i}(t)-x_{j}(t)\|^{2}=0$ (or $\lim_{t\to\infty}\|x_{i}(t)-x_{j}(t)\|=0$, a.s.).
Denote $x(t)=[x_{1}^{T}(t),...,x_{N}^{T}(t)]^{T}$. Substituting the protocol (2) into (1) yields the following closed-loop system
$\begin{split}dx(t)=&(I_{N}\otimes A+(\tilde{C}(t)\tilde{\Sigma}(t))\mathcal{L}\otimes B\mathcal{K})x(t)dt\\ &+(I_{N}\otimes C)x(t)dw(t),\end{split}$ (3)
where $\tilde{C}(t)=\mathrm{diag}\{c_{1}(t),...,c_{N}(t)\}$, and $\tilde{\Sigma}(t)=\mathrm{diag}\{\Sigma_{1}(t),...,\Sigma_{N}(t)\}$. Let $\vartheta_{i}(t)=x_{i}(t)-\frac{1}{N}\sum_{j=1}^{N}x_{j}(t)$. Denote $\vartheta(t)=[(I_{N}-\frac{1}{N}\mathbf{1}_{N}\mathbf{1}_{N}^{T})\otimes I_{n}]x(t)$, where $\vartheta(t)=[\vartheta_{1}^{T}(t),...,\vartheta_{N}^{T}(t)]^{T}$. Then, (3) can be written as
$\begin{split}d\vartheta(t)=&(I_{N}\otimes A+(\tilde{C}(t)\tilde{\Sigma}(t))\mathcal{L}\otimes B\mathcal{K})\vartheta(t)dt\\ &+(I_{N}\otimes C)\vartheta(t)dw(t).\end{split}$ (4)
The adaptive feedback law proposed in this paper introduces a historical integral term (e.g., $\int_{0}^{t}|r(s)x(s)|^{p}ds$ in [28]), resulting in a closed-loop system described by a path-dependent stochastic differential equation. Since the drift coefficient depends on the entire historical path of the solution, standard existence and uniqueness theorems under Lipschitz conditions no longer apply. Although deterministic counterparts can be treated with well-established methods, the stochastic path-dependent case remains theoretically underdeveloped. Therefore, as pointed in [28], it is necessary to establish the existence and uniqueness of solutions for this class of path-dependent stochastic systems.
III Existence and uniqueness of the solution to the path-dependent stochastic systems
Consider the following path-dependent stochastic system
$\begin{split}dy(t)\!=&\bigg[\!\int_{0}^{t}e^{\gamma s}y^{T}\!(s)Q_{1}y(s)ds)\bigg]\mathfrak{h}(y(t))dt+\mathfrak{f}(y(t))dt\\ &+\mathfrak{g}(y(t))dw(t),\end{split}$ (5)
where $y(t)\in\mathbb{R}^{n}$, $\gamma$ are constants, $Q_{1}\in\mathbb{R}^{n\times n}$ is constant matrices with $Q_{1}>0$, and $w(t)$ is a standard Brownian motion defined on the complete probability space $(\Omega,{\mathcal{F}},\mathbb{P})$. The nonlinear term $\mathfrak{f}$, $\mathfrak{g}$, $\mathfrak{h}$ satisfy the following assumption.
Assumption III.1 (Local Lipschitz Condition).
For each $q>0$ and $\forall y_{1},y_{2}\in\mathbb{R}^{n}$ with $\|y_{1}\|\vee\|y_{2}\|<q$, there exists a constant $L_{q}>0$ such that
$\begin{split}&\|\mathfrak{f}(y_{1})-\mathfrak{f}(y_{2})\|\vee\|\mathfrak{g}(y_{1})-\mathfrak{g}(y_{2})\|\vee\|\mathfrak{h}(y_{1})-\mathfrak{h}(y_{2})\|\\ &\leq L_{q}\|y_{1}-y_{2}\|.\end{split}$
In the interest of consistency, we introduce the following functional Itô formula [29]. Define $y_{t}=\{y(\theta):0\leq\theta\leq t\}$. Let $\overline{T}>0$ be a fixed terminal time. For each $t\in[0,\overline{T}]$, denote by $\Lambda_{t}$ the set of cadlag functions from $[0,t]$ to $\mathbb{R}^{n}$, and define $\Lambda\equiv\bigcup_{t\in[0,\overline{T}]}\Lambda_{t}$. Let $y$ be a continuous semimartingale process, with $y(t)$ its value at time $t$ and $y_{t}\in\Lambda_{t}$ its path over $[0,t]$. Let $U:\Lambda\times[0,\overline{T}]\to\mathbb{R}$ be a smooth functional in the sense of Dupire, i.e., $U$ is $\Lambda$-continuous, twice continuously differentiable in the spatial direction (in the sense of $\mathcal{\nabla}_{y}$) and once continuously differentiable in the temporal direction (in the sense of $\mathcal{\nabla}_{t}$), with these derivatives themselves $\Lambda$-continuous. Then the functional Itô formula (Theorem 3.1 of [29]) states that for any $t\in[0,\overline{T}]$:
$\begin{split}U(y_{t},t)&=U(y_{0},0)+\int_{0}^{t}\mathcal{\nabla}_{y}U(y_{s},s)\,dy(s)\\ &\quad+\int_{0}^{t}\mathcal{\nabla}_{t}U(y_{s},s)\,ds+\frac{1}{2}\int_{0}^{t}\mathcal{\nabla}_{yy}U(y_{s},s)\,d\langle y\rangle_{s},\end{split}$
where $\mathcal{\nabla}_{y}U(y_{t},t)$ and $\mathcal{\nabla}_{yy}U(y_{t},t)$ are the functional space derivatives, $\mathcal{\nabla}_{t}U(y_{t},t)$ is the functional time derivative, and $d\langle y\rangle_{t}$ is the quadratic variation of $y$. For precise definitions, see [29].
Based on the above functional Itô formula in [29], for path-dependent stochastic system (5) and a functional $U:\Lambda\times[0,\overline{T}]\to\mathbb{R}$, we have
$\begin{split}U(y_{t},t)=&U(y_{0},0)+\int_{0}^{t}\mathfrak{L}U(y_{s},s)ds\\ &+\int_{0}^{t}\mathcal{\nabla}_{y}U(y_{s},s)\mathfrak{g}(y(s))dw(s),\end{split}$
where $\mathfrak{L}U(y_{t},t)$ is defined as $\mathfrak{L}U(y_{t},t)=\mathcal{\nabla}_{t}U(y_{t},t)+\mathcal{\nabla}_{y}U(y_{t},t)\bar{\mathfrak{f}}(y_{t},t)+\frac{1}{2}{\rm{trace}}[\mathfrak{g}^{T}(y(t))\nabla_{yy}U(y_{t},t)\mathfrak{g}(y(t))]$ with $\bar{\mathfrak{f}}$ depends on the whole path $y_{t}$, $\bar{\mathfrak{f}}(y_{t},t)=\big[\int_{0}^{t}e^{\gamma s}y^{T}(s)Q_{1}y(s)ds)\big]\mathfrak{h}(y(t))+\mathfrak{f}(y(t))$.
To establish the existence and uniqueness of the solution to (5), the following lemma is introduced.
Lemma III.1.
Assume that there exist a functional $U:\Lambda\times[0,\overline{T}]\to\mathbb{R}$ and four constants $\alpha\geq 0$, $\beta\geq 0$, $k\leq 2$ and $\lambda_{1}\geq 0$ such that
(i) $e^{-\gamma t}\big(\!\!\int_{0}^{t}e^{\gamma s}y^{T}\!(s)Q_{1}y(s)ds-\beta\big)^{2}\vee\lambda_{1}|y(t)|^{2}\leq kU(y_{t},t)$;
(ii) $\mathfrak{L}U(y_{t},t)\!\leq\!\!-\alpha U(y_{t},t)\!+\phi(t)$,
where $\phi(t)\!=\!\frac{\alpha}{2}e^{-\gamma t}\big(\!\int_{0}^{t}e^{\gamma s}y^{T}\!(s)Q_{1}y(s)ds\!-\!\beta\big)^{2}$, $\gamma$ and $Q_{1}$ are defined in (5). Then the path-dependent stochastic system (5) exists a unique global solution for any initial value $y(0)\in\mathbb{R}^{n}$.
Proof.
By Assumption III.1, the drift and diffusion terms of the path-dependent stochastic system (5) satisfy the local Lipschitz condition. Therefore, for any initial value $y(0)\in\mathbb{R}^{n}$, there exists a unique maximal local solution $y(t)$ on $t\in[0,T_{\infty})$ to (5) due to the local Lipschitz continuity of the coefficient (see [28, 30]), where $T_{\infty}$ represents the explosion time. Let $h_{0}>0$. For each integer $h\geq h_{0}$, define the stopping time $\tau_{h}=\inf\{t\in[0,T_{\infty}):|y(t)|\geq h\}$. Denote $\tau_{\infty}=\lim_{h\to\infty}\tau_{h}$, where $\tau_{\infty}\leq T_{\infty}$ a.s. That is, if we can prove $\tau_{\infty}=\infty$ a.s., then $T_{\infty}=\infty$ follows, which means the existence and uniqueness of the solution. By the functional $\mathrm{It\hat{o}}$ formula and condition (ii), we have for any $h\geq h_{0}$
$\begin{split}&\mathbb{E}U(y_{\tau_{h}\wedge t},\tau_{h}\wedge t)-U(y_{0},0)\\ &\leq\mathbb{E}\int_{0}^{\tau_{h}\wedge t}\left(-\alpha U(y_{s},s)+\phi(s)\right)ds,\end{split}$ (6)
with $\phi(s)=\frac{\alpha}{2}e^{-\gamma s}\big(\int_{0}^{s}e^{\gamma r}y^{T}(r)Q_{1}y(r)ds-\beta\big)^{2}$. Note that $\phi(t)=\frac{\alpha}{2}e^{-\gamma t}\big(\int_{0}^{t}e^{\gamma s}y^{T}(s)Q_{1}y(s)ds-\beta\big)^{2}\leq\frac{\alpha k}{2}U(y_{t},t)$ from condition (i). Therefore, substituting this into (6) yields
$\begin{split}&\mathbb{E}U(y_{\tau_{h}\wedge t},\tau_{h}\wedge t)\\ &\leq U(y_{0},0)-\alpha(1-\frac{k}{2})\mathbb{E}\int_{0}^{\tau_{h}\wedge t}U(y_{s},s)ds\\ &\leq U(y_{0},0),\end{split}$ (7)
since $\alpha\geq 0$ and $k\leq 2$. Define $\theta_{h}=\inf_{|y|\geq h,t\geq 0}{U}(y)$, then
$\begin{split}\theta_{h}\mathbb{P}(\tau_{h}\leq t)\leq U(y_{0},0).\end{split}$ (8)
Since $\lim_{|y|\to\infty}\inf_{t\geq 0}{U}(y_{t},t)\geq\lim_{|y|\to\infty}\inf_{t\geq 0}\lambda_{1}|y|^{2}=\infty$ from condition (i), we have
$\begin{split}\theta_{h}\to\infty\ as\ h\to\infty.\end{split}$
Combining with (8), ones get
$\begin{split}\lim_{h\to\infty}\mathbb{P}(\tau_{h}\leq t)=0,\end{split}$
which implies, $\mathbb{P}(\tau_{\infty}\leq t)=0$, namely $\mathbb{P}(\tau_{\infty}>t)=1$. Since $t\geq 0$ is arbitrary, $\mathbb{P}(\tau_{\infty}=\infty)=1$. Therefore, the existence and uniqueness of the solution can be proved. ∎
Remark III.1.
Lemma III.1 established a sufficient condition for proving the existence of a global solution to the path-dependent stochastic system (5). Its core mechanism relies on constructing a Lyapunov functional $U(y_{t},t)$ satisfying the inequality $\mathfrak{L}U(y_{t},t)\leq-\alpha U(y_{t},t)+\phi(t)$, where the path-dependent term $\phi(t)$ is instantaneously and pointwise bounded by $U(y_{t},t)$ via the coupling condition (i). While this condition is pivotal to the proof, it constitutes the primary restriction, as it requires the path-dependent term $\phi(t)$ to be strictly scaled by the current state energy $U(y_{t},t)$ at every moment.
Based on this lemma, the subsequent consensus analysis in this paper is conducted under the premise of the existence and uniqueness of the solution, thereby ensuring the rigor of the theoretical analysis.
IV Fully distributed consensus control for stochastic MASs
In this section, the unified fully distributed protocol (2) is specifically designed to solve the consensus problem of stochastic MASs for the cases of undirected and directed graphs, respectively.
As we all know that the algebraic Riccati equation is an important tool for the control design of linear deterministic systems [31, 32]. For linear stochastic systems, we introduce the following stochastic algebraic Riccati equation(SARE) to design the controller parameters
$\begin{split}A^{T}P+PA-PBB^{T}P+C^{T}PC+I_{n}=0.\end{split}$ (9)
Note that the solvability of this SARE has been well investigated in [33], which demonstrates that the existence and uniqueness of the positive definite solution $P$ is equivalent to that the corresponding stochastic system $dx(t)=[Ax(t)+Bu(t)]dt+Cx(t)dw(t)$ is m.s. stabilizable. So in this work, SARE (9) is always assumed to have a positive-definite solution $P$.
IV-A Fully distributed consensus control for stochastic MASs with the directed topology
Lemma IV.1 ([34]).
Let $\tilde{\mathcal{G}}$ be a directed graph that contains a spanning tree. Then the node set $\mathcal{V}$ can be partitioned into two subsets $\mathcal{V}_{l}$ and $\mathcal{V}_{f}$ satisfying the following properties:
  1. (i)
    $\mathcal{V}_{l}\cup\mathcal{V}_{f}=\mathcal{V}$ and $\mathcal{V}_{l}\cap\mathcal{V}_{f}=\emptyset$;
  2. (ii)
    the subgraph induced by $\mathcal{V}_{l}$ is strongly connected;
  3. (iii)
    there is no edge from any node in $\mathcal{V}_{l}$ to any node in $\mathcal{V}_{f}$.
By relabeling the nodes such that $\mathcal{V}_{l}=\{n_{1},\dots,n_{M}\}$ and $\mathcal{V}_{f}=\{n_{M+1},\dots,n_{N}\}$ with $1\leq M\leq N$, the Laplacian matrix $\mathcal{L}\in\mathbb{R}^{N\times N}$ of $\tilde{\mathcal{G}}$ can be written in the following block form:
$\mathcal{L}=\begin{bmatrix}\mathcal{L}_{11}&0\\ \mathcal{L}_{21}&\mathcal{L}_{22}\end{bmatrix}$,
where $\mathcal{L}_{11}\in\mathbb{R}^{M\times M}$ is the Laplacian matrix of the strongly connected subgraph induced by $\mathcal{V}_{l}$, $\mathcal{L}_{22}\in\mathbb{R}^{(N-M)\times(N-M)}$ is a nonsingular $M$-matrix, and $\mathcal{L}_{21}$ represents the connectivity between $\mathcal{V}_{l}$ and $\mathcal{V}_{f}$.
Lemma IV.2 ([12]).
Let $\mathcal{L}_{11}\in\mathbb{R}^{M\times M}$ with $M\geq 1$ be the Laplacian matrix of the strongly connected subgraph induced by $\mathcal{V}_{l}$. Then there exists a positive diagonal matrix $R=\operatorname{diag}\{r_{1},\dots,r_{M}\}$ such that $\tilde{\mathcal{L}}_{11}=R\mathcal{L}_{11}+\mathcal{L}_{11}^{T}R$ is symmetric positive semidefinite. Moreover, its eigenvalues can be written as $0=\lambda_{1}(\tilde{\mathcal{L}}_{11})<\lambda_{2}(\tilde{\mathcal{L}}_{11})\leq\cdots\leq\lambda_{M}(\tilde{\mathcal{L}}_{11})$. In particular, we can choose $R$ as $r=[r_{1},\dots,r_{M}]^{T}$ with $\sum_{i=1}^{M}r_{i}=1$ and $r^{T}\mathcal{L}_{11}=0$.
Lemma IV.3 ([35]).
Let $\mathcal{L}_{22}\in\mathbb{R}^{(N-M)\times(N-M)}$ with $M\leq N$ be the nonsingular $M$-matrix. Then there exists a positive diagonal matrix $S=\operatorname{diag}\{s_{1},\dots,s_{N-M}\}$ such that $\tilde{\mathcal{L}}_{22}=S\mathcal{L}_{22}+\mathcal{L}_{22}^{T}S$ is symmetric positive definite. Consequently, its eigenvalues satisfy $0<\lambda_{1}(\tilde{\mathcal{L}}_{22})\leq\lambda_{2}(\tilde{\mathcal{L}}_{22})\leq\cdots\leq\lambda_{N-M}(\tilde{\mathcal{L}}_{22})$. A particular choice of matrix $S$ is $s=[s_{1},\dots,s_{N-M}]^{T}=(\mathcal{L}_{22}^{T})^{-1}\mathbf{1}_{N-M}$.
Let $\xi(t)=[\xi_{1}^{T}(t),...,\xi_{N}^{T}(t)]^{T}$ and $\tilde{\sigma}(t)=\mathrm{diag}\{\sigma_{1}(t),...,\sigma_{N}(t)\}$. According to Lemma IV.1, the states $\xi(t)$ can be partitioned into $\xi_{l}(t)=[\xi_{1}^{T}(t),\dots,\xi_{M}^{T}(t)]^{T}$ and $\xi_{f}(t)=[\xi_{M+1}^{T}(t),\dots,\xi_{N}^{T}(t)]^{T}$, the states $\vartheta(t)$ can be partitioned into $\vartheta_{l}(t)=[\vartheta_{1}^{T}(t),\dots,\vartheta_{M}^{T}(t)]^{T}$ and $\vartheta_{f}(t)=[\vartheta_{M+1}^{T}(t),\dots,\vartheta_{N}^{T}(t)]^{T}$, the time-varying adaptive gain $\tilde{C}(t)$ can be partitioned into $\tilde{C}_{l}(t)=\mathrm{diag}\{c_{1}(t),...,c_{M}(t)\}$ and $\tilde{C}_{f}(t)=\mathrm{diag}\{c_{M+1}(t),...,c_{N}(t)\}$, the auxiliary time-varying gain $\tilde{\Sigma}(t)$ can be partitioned into $\tilde{\Sigma}_{l}(t)=\mathrm{diag}\{\Sigma_{1}(t),...,\Sigma_{M}(t)\}$ and $\tilde{\Sigma}_{f}(t)=\mathrm{diag}\{\Sigma_{M+1}(t),...,\Sigma_{N}(t)\}$, $\tilde{\sigma}(t)$ can be partitioned into $\tilde{\sigma}_{l}(t)=\mathrm{diag}\{\sigma_{1}(t),...,\sigma_{M}(t)\}$ and $\tilde{\sigma}_{f}(t)=\mathrm{diag}\{\sigma_{M+1}(t),...,\sigma_{N}(t)\}$, corresponding to the two subsets $\mathcal{V}_{l}$ and $\mathcal{V}_{f}$. Based on the closed-loop stochastic system (4), we obtain $d\vartheta_{l}(t)=(I_{M}\otimes A+(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t)\mathcal{L}_{11})\otimes B\mathcal{K})\vartheta_{l}(t)dt+(I_{M}\otimes C)\vartheta_{l}(t)dw(t)$, and $d\vartheta_{f}(t)=(I_{N-M}\otimes A+(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t)\mathcal{L}_{22})\otimes B\mathcal{K})\vartheta_{f}(t)dt+(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t)\mathcal{L}_{21}\otimes B\mathcal{K})\vartheta_{l}(t)dt+(I_{N-M}\otimes C)\vartheta_{f}(t)dw(t)$. Note that $\xi_{l}(t)=(\mathcal{L}_{11}\otimes I_{n})\vartheta_{l}(t)$ and $\xi_{f}(t)=(\mathcal{L}_{21}\otimes I_{n})\vartheta_{l}(t)+(\mathcal{L}_{22}\otimes I_{n})\vartheta_{f}(t)$. Then, we can get
$\begin{split}d\xi_{l}(t)=&(I_{M}\otimes A+(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t)\mathcal{L}_{11})\otimes B\mathcal{K})\xi_{l}(t)dt\\ &+(I_{M}\otimes C)\xi_{l}(t)dw(t),\end{split}$ (10)
and
$\begin{split}d\xi_{f}(t)=&(I_{N-M}\otimes A+(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t)\mathcal{L}_{22})\otimes B\mathcal{K})\xi_{f}(t)dt\\ &+(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t)\mathcal{L}_{21}\otimes B\mathcal{K})\xi_{l}(t)dt\\ &+(I_{N-M}\otimes C)\xi_{f}(t)dw(t).\end{split}$ (11)
For the case of $\mu>1$ and $\gamma=0$, we have the following theorem.
Theorem IV.1.
Suppose the directed graph $\tilde{\mathcal{G}}$ contains a spanning tree. The m.s. and a.s. consensus of (1) can be solved by the protocol (2) with $\gamma=0$, $\mathcal{\mathcal{K}}=-B^{T}P$, and $\Gamma=PBB^{T}P$, where $P$ is the solution to the SARE (9). The auxiliary time-varying gain $\Sigma_{i}(t)$ satisfies
(I) $\Sigma_{i}(t)=k_{1}(k_{2}+\frac{\sigma_{i}(t)}{c_{i}(t)})^{\mu}$, with $\mu>1,k_{1},k_{2}\geq 1$;
(II) $c_{i}(0)\geq 1$, $\Sigma_{i}(0)\geq 1$.
Proof.
Based on the closed-loop stochastic system (4), (10) and (11), we choose the following Lyapunov functional
$V(\vartheta_{t},t)=V_{l}(\vartheta_{t},t)+V_{f}(\vartheta_{t},t)$
with $V_{l}(\vartheta_{t},t)=\eta_{1}^{l}\sum_{i=1}^{M}r_{i}c_{i}(t)\int_{0}^{\sigma_{i}(t)}k_{1}(k_{2}+\frac{s}{c_{i}(t)})^{\mu}ds+\eta_{2}^{l}\sum_{i=1}^{M}r_{i}(c_{i}(t)-\psi_{i})^{2}$ and $V_{f}(\vartheta_{t},t)=\eta_{1}^{f}\sum_{i=M+1}^{N}s_{i-M}c_{i}(t)\int_{0}^{\sigma_{i}(t)}k_{1}(k_{2}+\frac{s}{c_{i}(t)})^{\mu}ds+\eta_{2}^{f}\sum_{i=M+1}^{N}s_{i-M}(c_{i}(t)-\psi_{i})^{2}$, where $\eta_{1}^{l},\eta_{2}^{l},\eta_{1}^{f},\eta_{2}^{f}\geq 0$, $r_{i},s_{i}>0$, and $\psi_{i}$ are positive constants to be determined. According to the functional $\mathrm{It\hat{o}}$ formula and condition (I), it can be deduced that $dV(\vartheta_{t},t)$ has the following form
$\begin{split}dV&(\vartheta_{t},t)\!=\!\mathfrak{L}V(\vartheta_{t},t)dt+\frac{\partial V(\vartheta_{t},t)}{\partial\vartheta}(I_{N}\!\otimes\!C)\vartheta(t)dw(t),\end{split}$
where $\mathfrak{L}V(\vartheta_{t},t)$ is defined as
$\begin{split}\mathfrak{L}V(\vartheta_{t},t)=\mathfrak{L}V_{l}(\vartheta_{t},t)+\mathfrak{L}V_{f}(\vartheta_{t},t)\end{split}$ (12)
with $\mathfrak{L}V_{l}(\vartheta_{t},t)=\eta_{1}^{l}\sum_{i=1}^{M}r_{i}\dot{c}_{i}(t)\int_{0}^{\sigma_{i}(t)}k_{1}(k_{2}+\frac{s}{c_{i}(t)})^{\mu}ds+2\eta_{2}^{l}\sum_{i=1}^{M}r_{i}(c_{i}(t)-\psi_{i})\dot{c}_{i}(t)+\eta_{1}^{l}\sum_{i=1}^{M}r_{i}c_{i}(t)\Sigma_{i}(t)\mathfrak{L}\sigma_{i}(t)-\int_{0}^{\sigma_{i}(t)}\mu k_{1}\frac{s}{c_{i}^{2}(t)}\dot{c}_{i}(t)(k_{2}+\frac{s}{{c}_{i}(t)})^{\mu-1}ds$, and $\mathfrak{L}V_{f}(\vartheta_{t},t)=\eta_{1}^{f}\sum_{i=M+1}^{N}s_{i-M}\dot{c}_{i}(t)\int_{0}^{\sigma_{i}(t)}k_{1}(k_{2}+\frac{s}{c_{i}(t)})^{\mu}ds+2\eta_{2}^{f}\sum_{i=M+1}^{N}s_{i-M}(c_{i}(t)-\psi_{i})\dot{c}_{i}(t)+\eta_{1}^{f}\sum_{i=M+1}^{N}s_{i-M}c_{i}(t)\Sigma_{i}(t)\mathfrak{L}\sigma_{i}(t)-\int_{0}^{\sigma_{i}(t)}\mu k_{1}\frac{s}{c_{i}^{2}(t)}\dot{c}_{i}(t)(k_{2}+\frac{s}{{c}_{i}(t)})^{\mu-1}ds$. Since $\mu,k_{1},k_{2}>0,\sigma_{i}(t)\geq 0$, we have $\mathfrak{L}V_{l}(\vartheta_{t},t)\leq\sum_{i=1}^{M}r_{i}\big[\eta_{1}^{l}\Sigma_{i}(t)\sigma_{i}(t)+2\eta_{2}^{l}c_{i}(t)-2\eta_{2}^{l}\psi_{i}\big]\dot{c}_{i}(t)+\eta_{1}^{l}\sum_{i=1}^{M}r_{i}c_{i}(t)\Sigma_{i}(t)\mathfrak{L}\sigma_{i}(t)$, and $\mathfrak{L}V_{f}(\vartheta_{t},t)\leq\sum_{i=M+1}^{N}s_{i-M}\big[\eta_{1}^{f}\Sigma_{i}(t)\sigma_{i}(t)+2\eta_{2}^{f}c_{i}(t)-2\eta_{2}^{f}\psi_{i}\big]\dot{c}_{i}(t)+\eta_{1}^{f}\sum_{i=M+1}^{N}s_{i-M}c_{i}(t)\Sigma_{i}(t)\mathfrak{L}\sigma_{i}(t)$. Let $\tilde{\psi}=\mathrm{diag}\{\psi_{1},...,\psi_{N}\}$. According to Lemma IV.1, $\tilde{\psi}$ can be partitioned into $\tilde{\psi}_{l}=\mathrm{diag}\{\psi_{1},...,\psi_{M}\}$ and $\tilde{\psi}_{f}=\mathrm{diag}\{\psi_{M+1},...,\psi_{N}\}$. Let $\mathcal{K}=-B^{T}P$ and $\Gamma=PBB^{T}P$. Note that $\sum_{i=1}^{M}\sigma_{i}(t)=\xi_{l}^{T}(t)(I_{M}\otimes P)\xi_{l}(t)$ and $\xi_{l}(t)=(\mathcal{L}_{11}\otimes I_{n})\vartheta_{l}(t)$. By the definition of $\sigma_{i}(t)$, (10), (11) and using Lemma IV.1 and IV.2, we have
$\begin{split}&\sum_{i=1}^{M}r_{i}c_{i}(t)\Sigma_{i}(t)\mathfrak{L}\sigma_{i}(t)\\ &=\xi_{l}^{T}\!(t)[R(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\xi_{l}(t)\!\!\!\!\\ &\ \ \ -\!\xi_{l}^{T}(t)[(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))^{2}\tilde{\mathcal{L}}_{11}\otimes\Gamma]\xi_{l}(t),\end{split}$ (13)
and
$\begin{split}&\sum_{i=1}^{M}r_{i}[\eta_{1}^{l}\Sigma_{i}(t)\sigma_{i}(t)+2\eta_{2}^{l}c_{i}(t)-2\eta_{2}^{l}\psi_{i}]\dot{c}_{i}(t)\\ &=\xi_{l}^{T}\!(t)[R(\eta_{1}^{l}\tilde{\Sigma}_{l}(t)\tilde{\sigma}_{l}(t)+2\eta_{2}^{l}(\tilde{C}_{l}(t)\!-\!\tilde{\psi}_{l}))\!\otimes\!\Gamma]\xi_{l}(t),\end{split}$ (14)
Note that $\sum_{i=M+1}^{N}\sigma_{i}(t)=\xi_{f}^{T}(t)(I_{N-M}\otimes P)\xi_{f}(t)$. Then, using Lemma IV.1, IV.3, and the elementary inequality: $2a^{T}Ob\leq\iota a^{T}Oa+\frac{1}{\iota}b^{T}Ob$, for any positive definite matrix $O\in\mathbb{R}^{n\times n}$, $a,b\in\mathbb{R}^{n}$ and $\iota>0$, we have
$\begin{split}&\sum_{i=M+1}^{N}s_{i-M}c_{i}(t)\Sigma_{i}(t)\mathfrak{L}\sigma_{i}(t)\\ &=\xi_{f}^{T}\!(t)[S(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\xi_{f}(t)\!\!\!\!\\ &\ \ \ -\!\xi_{f}^{T}(t)[(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))^{2}\tilde{\mathcal{L}}_{22}\otimes\Gamma]\xi_{f}(t)\\ &\ \ \ -\!2\xi_{f}^{T}\!(t)[(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))S\mathcal{L}_{21}\!\otimes\!\Gamma]\xi_{l}(t)\\ &\leq\xi_{f}^{T}\!(t)[S(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\xi_{f}(t)\!\!\!\!\\ &\ \ \ -\xi_{f}^{T}(t)[(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))^{2}\tilde{\mathcal{L}}_{22}\otimes\Gamma]\xi_{f}(t)\\ &\ \ \ +\iota\xi_{f}^{T}\!(t)[(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))^{2}\!\otimes\!\Gamma]\xi_{f}(t)\\ &\ \ \ +\frac{\sigma_{\max}^{2}(S\mathcal{L}_{21})}{\iota}\xi_{l}^{T}\!(t)[(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))^{2}\!\otimes\!\Gamma]\xi_{l}(t),\end{split}$ (15)
and
$\begin{split}&\sum_{i=M+1}^{N}s_{i-M}[\eta_{1}^{f}\Sigma_{i}(t)\sigma_{i}(t)+2\eta_{2}^{f}c_{i}(t)-2\eta_{2}^{f}\psi_{i}]\dot{c}_{i}(t)\\ &=\xi_{f}^{T}\!(t)[S(\eta_{1}^{f}\tilde{\Sigma}_{f}(t)\tilde{\sigma}_{f}(t)\!+\!2\eta_{2}^{f}(\tilde{C}_{f}(t)\!-\!\tilde{\psi}_{f}))\!\otimes\!\Gamma]\xi_{f}(t),\end{split}$ (16)
By substituting (13)-(16) into (12), we can obtain
$\begin{split}\mathfrak{L}&V(\vartheta_{t},t)\\ \leq&\eta_{1}^{l}\xi_{l}^{T}\!(t)[R(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))\otimes(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\xi_{l}(t)\\ &\!\!\!\!\!+\!\eta_{1}^{f}\xi_{f}^{T}\!(t)[S(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\xi_{f}(t)\\ &+\xi_{l}^{T}\!(t)\!\Big[\Phi_{t}^{l}\otimes\!\Gamma\Big]\!\xi_{l}(t)+\xi_{f}^{T}\!(t)\!\Big[\Phi_{t}^{f}\otimes\!\Gamma\Big]\!\xi_{f}(t),\end{split}$ (17)
with $\Phi_{t}^{l}=\eta_{1}^{l}R\tilde{\Sigma}_{l}(t)\tilde{\sigma}_{l}(t)+2\eta_{2}^{l}R(\tilde{C}_{l}(t)-\tilde{\psi}_{l})+\eta_{1}^{f}\frac{\sigma_{\max}^{2}(S\mathcal{L}_{21})}{\iota}(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))^{2}-\eta_{1}^{l}\tilde{\mathcal{L}}_{11}(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))^{2}$ and $\Phi_{t}^{f}=\eta_{1}^{f}S\tilde{\Sigma}_{f}(t)\tilde{\sigma}_{f}(t)+2\eta_{2}^{f}S(\tilde{C}_{f}(t)-\tilde{\psi}_{f})+\eta_{1}^{f}\iota(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))^{2}-\eta_{1}^{f}\tilde{\mathcal{L}}_{22}(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))^{2}$. Note that $\Sigma_{i}(t)=k_{1}(k_{2}+\frac{\sigma_{i}(t)}{c_{i}(t)})^{\mu}$, for $\mu>1,k_{1},k_{2}\geq 1$ and $c_{i}(t)\geq 1$, $\Sigma_{i}(t)\geq 1$ by condition (II). Therefore, we have $\tilde{\sigma}_{l}(t)\leq\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t)^{\frac{1}{\mu}}$ and $\tilde{\sigma}_{f}(t)\leq\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t)^{\frac{1}{\mu}}$. Together with the Young’s inequality: $ab\leq\epsilon a^{p}+\frac{(\epsilon p)^{-q/p}}{q}b^{q}$ with $p,q>1$ and $\frac{1}{p}+\frac{1}{q}=1$ for any $\epsilon>0$ and $a,b\geq 0$, we have
$\begin{split}\Phi_{t}^{l}\leq&\eta_{1}^{l}R(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))^{1+\frac{1}{\mu}}+2\eta_{2}^{l}R(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t)-\tilde{\psi}_{l})\\ &-\Big(\eta_{1}^{l}\tilde{\mathcal{L}}_{11}-\eta_{1}^{f}\frac{\sigma_{\max}^{2}(S\mathcal{L}_{21})}{\iota}I_{M}\Big)(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))^{2}\\ \leq&\eta_{1}^{l}\Big(\epsilon_{1}[(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))^{1+\frac{1}{\mu}}]^{p_{1}}+\frac{(\epsilon_{1}p_{1})^{-q_{1}/p_{1}}}{q_{1}}R^{q_{1}}\Big)\\ &+2\eta_{2}^{l}\Big(\epsilon_{2}(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))^{p_{2}}+\frac{(\epsilon_{2}p_{2})^{-q_{2}/p_{2}}}{q_{2}}R^{q_{2}}\Big)\\ &-\!2\eta_{2}^{l}R\tilde{\psi}_{l}\!-\!\!\Big(\!\eta_{1}^{l}\lambda_{2}(\tilde{\mathcal{L}}_{11})\!-\!\eta_{1}^{f}\frac{\sigma_{\max}^{2}(S\mathcal{L}_{21})}{\iota}\!\Big)\!(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))^{2}\!.\end{split}$
Note that we can choose $p_{1}$ and $p_{2}$ to satisfy $(1+\frac{1}{\mu})p_{1}\leq 2$ and $1<p_{2}\leq 2$ since $\mu>1$ and $p_{1},p_{2}>1$. Combining these conditions yields
$\begin{split}\Phi_{t}^{l}\!\leq&\Big[\!\eta_{1}^{l}\epsilon_{1}\!+\!2\eta_{2}^{l}\epsilon_{2}\!-\!\!\Big(\!\eta_{1}^{l}\!\lambda_{2}(\tilde{\mathcal{L}}_{11})\!-\!\eta_{1}^{f}\frac{\sigma_{\max}^{2}\!(\!S\mathcal{L}_{21}\!)}{\iota}\!\Big)\!\Big]\!(\!\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t)\!)^{2}\\ &-\!2\eta_{2}^{l}R\tilde{\psi}_{l}\!+\!\eta_{1}^{l}\frac{(\epsilon_{1}p_{1})^{-q_{1}/p_{1}}}{q_{1}}R^{q_{1}}\!+\!2\eta_{2}^{l}\frac{(\epsilon_{2}p_{2})^{-q_{2}/p_{2}}}{q_{2}}R^{q_{2}}.\end{split}$
Let $\psi_{i}\!\geq\!\frac{\eta_{1}^{l}(\epsilon_{1}p_{1})^{-q_{1}/p_{1}}}{2\eta_{2}^{l}q_{1}}\{r_{i}\}^{q_{1}-1}\!+\frac{(\epsilon_{2}p_{2})^{-q_{2}/p_{2}}}{q_{2}}\{r_{i}\}^{q_{2}-1}+\frac{\check{\psi_{i}}}{2\eta_{2}^{l}r_{i}}$ for $i=1,\dots,M$, where $\check{\psi_{i}}$, $i=1,\dots,M$ will be designed later. Therefore, we obtain
$\begin{split}\Phi_{t}^{l}&\leq-\eta_{3}^{l}(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))^{2}-\check{\psi}_{l}.\end{split}$ (18)
with $\eta_{3}^{l}=\eta_{1}^{l}\lambda_{2}(\tilde{\mathcal{L}}_{11})-\eta_{1}^{f}\frac{\sigma_{\max}^{2}(S\mathcal{L}_{21})}{\iota}-\eta_{1}^{l}\epsilon_{1}-2\eta_{2}^{l}\epsilon_{2}$ and $\check{\psi}_{l}=\mathrm{diag}\{\check{\psi}_{1},...,\check{\psi}_{M}\}$. Let $\psi_{i}\!\geq\!\frac{\eta_{1}(\epsilon_{1}p_{1})^{-q_{1}/p_{1}}}{2\eta_{2}q_{1}}\{s_{i}\}^{q_{1}-1}\!+\frac{(\epsilon_{2}p_{2})^{-q_{2}/p_{2}}}{q_{2}}\{s_{i}\}^{q_{2}-1}+\frac{\check{\psi_{i}}}{2\eta_{2}s_{i}}$ for $i=M+1,\dots,N$, where $\check{\psi_{i}}$, $i=M+1,\dots,N$ will be designed later. Similarly, we have
$\begin{split}\Phi_{t}^{f}&\leq-\eta_{3}^{f}(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))^{2}-\check{\psi}_{f}.\end{split}$ (19)
with $\eta_{3}^{f}=\eta_{1}^{f}(\lambda_{1}(\tilde{\mathcal{L}}_{22})-\iota)-\eta_{1}^{f}\epsilon_{1}-2\eta_{2}^{f}\epsilon_{2}$ and $\check{\psi}_{f}=\mathrm{diag}\{\check{\psi}_{M+1},...,\check{\psi}_{N}\}$. Choose $\eta_{1}^{f}<\frac{\eta_{1}^{l}\lambda_{1}(\tilde{\mathcal{L}}_{22})\lambda_{2}(\tilde{\mathcal{L}}_{11})}{\sigma_{\max}^{2}(S\mathcal{L}_{21})}$, $\iota<\lambda_{1}(\tilde{\mathcal{L}}_{22})$, and sufficiently small $\epsilon_{1},\epsilon_{2}$ to ensure $\eta_{3}^{l}>0$ and $\eta_{3}^{f}>0$. Substituting the above inequalities (18) and (19) into (17) and using the elementary inequality: $2a^{T}Ob\leq\iota a^{T}Oa+\frac{1}{\iota}b^{T}Ob$, for any positive definite matrix $O\in\mathbb{R}^{n\times n}$, $a,b\in\mathbb{R}^{n}$ and $\iota>0$, ones get
$\begin{split}\mathfrak{L}&V(\vartheta_{t},t)\\ \leq&\eta_{1}^{l}\xi_{l}^{T}\!(t)[R(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))\otimes(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\xi_{l}(t)\\ &+\!\eta_{1}^{f}\xi_{f}^{T}\!(t)[S(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\xi_{f}(t)\\ &-\xi_{l}^{T}\!(t)\Big[\big(\eta_{3}^{l}(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))^{2}\!+\!\check{\psi}_{l}\big)\otimes\Gamma\Big]\xi_{l}(t)\\ &-\xi_{f}^{T}\!(t)\Big[\big(\eta_{3}^{f}(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))^{2}\!+\!\check{\psi}_{f}\big)\otimes\Gamma\Big]\xi_{f}(t),\\ \leq&\eta_{1}^{l}\xi_{l}^{T}\!(t)[R(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))\otimes(A^{T}P\!+\!PA\!+\!C^{T}\!PC)]\xi_{l}(t)\\ &-\xi_{l}^{T}(t)\left[2\sqrt{\eta_{3}^{l}\check{\psi}_{l}}(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))\otimes\Gamma\right]\xi_{l}(t)\\ &+\eta_{1}^{f}\xi_{f}^{T}\!(t)[S(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))\otimes(A^{T}P\!+\!PA\!+\!C^{T}\!PC)]\xi_{f}(t)\\ &-\xi_{f}^{T}(t)\left[2\sqrt{\eta_{3}^{f}\check{\psi}_{f}}(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))\otimes\Gamma\right]\xi_{f}(t).\end{split}$
Let $\check{\psi_{i}}\geq\frac{(\eta_{1}^{l}\max\{r_{i}\})^{2}}{4\eta_{3}^{l}}$ for $i=1,\dots,M$ and $\check{\psi_{i}}\geq\frac{(\eta_{1}^{f}\max\{s_{i-M}\})^{2}}{4\eta_{3}^{f}}$ for $i=M+1,\dots,N$. Then, we have
$\begin{split}\mathfrak{L}V(\vartheta_{t},t)\leq&-\eta_{1}^{l}\xi_{l}^{T}(t)[R(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))\otimes I_{n}]\xi_{l}(t)\\ &-\eta_{1}^{f}\xi_{f}^{T}(t)[S(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))\otimes I_{n}]\xi_{f}(t)\\ \leq&-\eta_{1}^{l}\varepsilon\xi_{l}^{T}(t)[R(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))\otimes P]\xi_{l}(t)\\ &-\eta_{1}^{f}\varepsilon\xi_{f}^{T}(t)[S(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))\otimes P]\xi_{f}(t)\\ =&-\eta_{1}^{l}\varepsilon\sum_{i=1}^{M}r_{i}c_{i}(t)\Sigma_{i}(t)\sigma_{i}(t)\\ &-\eta_{1}^{f}\varepsilon\sum_{i=M+1}^{N}s_{i-M}c_{i}(t)\Sigma_{i}(t)\sigma_{i}(t)\\ \leq&-\eta_{1}^{l}\varepsilon\sum_{i=1}^{M}r_{i}c_{i}(t)\!\!\int_{0}^{\sigma_{i}(t)}\!\!\!k_{1}\Big(k_{2}+\frac{s}{c_{i}(t)}\Big)^{\mu}ds\\ &-\eta_{1}^{f}\varepsilon\!\!\!\sum_{i=M+1}^{N}\!\!\!\!\!s_{i-M}c_{i}(t)\!\!\!\int_{0}^{\sigma_{i}(t)}\!\!\!\!\!\!k_{1}\Big(k_{2}\!+\!\frac{s}{c_{i}(t)}\Big)^{\mu}ds\\ \leq&-\varepsilon V(\vartheta_{t},t)+\varphi(t),\end{split}$ (20)
where $A^{T}P+PA-PBB^{T}P+C^{T}PC+I_{n}=0$, $\varepsilon=\frac{1}{\lambda_{\max}(P)}$, and $\varphi(t)=\eta_{2}^{l}\varepsilon\sum_{i=1}^{M}r_{i}(c_{i}(t)-\psi_{i})^{2}+\eta_{2}^{f}\varepsilon\sum_{i=M+1}^{N}s_{i-M}(c_{i}(t)-\psi_{i})^{2}$. Next, we will prove the existence and uniqueness, as well as m.s. and a.s. stability of the solution to (4) in the following two steps.
Step 1: We need to prove the existence and uniqueness of the solution to the closed-loop stochastic system (4), as the system is inherently highly nonlinear and no longer satisfy the linear growth condition. Since $(I_{N}\otimes A+(\tilde{C}(t)\tilde{\Sigma}(t))\mathcal{L}\otimes B\mathcal{K})\vartheta(t)\leq(I_{N}\otimes A)\vartheta(t)+k_{1}\big[\big(\int_{0}^{t}\vartheta^{T}(s)(\mathcal{L}^{T}\mathcal{L}\otimes PBB^{T}P)\vartheta(s)ds+C_{1}(0)\big)\big(k_{2}+\vartheta^{T}(t)(\mathcal{L}^{T}\mathcal{L}\otimes P)\vartheta(t)\big)^{\mu}\big]\mathcal{L}\otimes B\mathcal{K}]\vartheta(t)$ where $C_{1}(0)\geq 0$ holds, the drift and diffusion terms of closed-loop stochastic system (4) satisfy the local Lipschitz condition. According to Lemma III.1, the existence and uniqueness of the solution to (4) can be proved.
Step 2: By the functional $\mathrm{It\hat{o}}$ formula for $V(\vartheta_{t},t)$ and the estimation (20) of $\mathfrak{L}V(\vartheta_{t},t)\leq-\eta_{1}^{l}\varepsilon\xi_{l}^{T}(t)[R(\tilde{C}_{l}(t)\tilde{\Sigma}_{l}(t))\otimes P]\xi_{l}(t)-\eta_{1}^{f}\varepsilon\xi_{f}^{T}(t)[S(\tilde{C}_{f}(t)\tilde{\Sigma}_{f}(t))\otimes P]\xi_{f}(t)\leq-\eta_{1}\varepsilon\xi^{T}(t)[\tilde{S}(\tilde{C}(t)\tilde{\Sigma}(t))\otimes P]\xi(t)$ with $\eta_{1}=\eta_{1}^{l}\wedge\eta_{1}^{f}$ and $\tilde{S}=\mathrm{diag}\{R,S\}$, it can be deduced that
$\begin{split}&V(\vartheta_{t},t)-V(\vartheta_{0},0)\\ &=\int_{0}^{t}\mathfrak{L}V(\vartheta_{s},s)ds+M_{v1}(t)\\ &\leq-\eta_{1}\varepsilon\!\int_{0}^{t}\!\xi^{T}(s)[\tilde{S}(\tilde{C}(s)\tilde{\Sigma}(s))\!\otimes\!P]\xi(s)ds+M_{v1}(t),\end{split}$ (21)
where $M_{v1}(t)$ is a local martingale with $M_{v1}(0)=0$. Using the non-negative semi-martingale convergence theorem, ones have
$\begin{split}\lim_{t\to\infty}\sup V(\vartheta_{t},t)<\infty\ a.s.\end{split}$
In addition, taking the expectations on both sides of (21), we can obtain
$\begin{split}\ \mathrm{E}\!\!\int_{0}^{t}\!\!\eta_{1}\varepsilon\xi^{T}\!(s)[\tilde{S}(\tilde{C}(s)\tilde{\Sigma}(s))\!\otimes\!P]\xi(s)ds\leq\!\mathrm{E}V(\vartheta_{0},0)\!<\!\infty,\end{split}$
that is, letting $t\to\infty$ yields
$\begin{split}\mathrm{E}\int_{0}^{\infty}\eta_{1}\varepsilon\xi^{T}(s)[\tilde{S}(\tilde{C}(s)\tilde{\Sigma}(s))\otimes P]\xi(s)ds<\infty.\end{split}$
Combining the above inequality with Barbalat’s lemma [36], we have $\lim_{t\to\infty}\mathrm{E}\xi^{T}(t)[\eta_{1}\varepsilon\tilde{S}(\tilde{C}(t)\tilde{\Sigma}(t))\otimes P]\xi(t)=0$, that is, $\lim_{t\to\infty}\mathrm{E}|(\tilde{C}(t)\tilde{\Sigma}(t))|\|\xi(t)\|^{2}=0$. Note that $\xi(t)=(\mathcal{L}\otimes I_{n})\vartheta(t)$ and $\lim_{|\vartheta|\to\infty}\inf_{t\geq 0}V(\vartheta_{t},t)=\infty$. Using the stochastic LaSalle lemma [37, Corollary 3.1], we can get $\lim_{t\to\infty}\xi(t)=0$ a.s. According to the definition of $\xi(t)$, the m.s. and a.s. consensus can be solved. ∎
Remark IV.1.
In fact, the function $\Sigma_{i}$ admits other design forms. Especially, when we choose $\Sigma_{i}(t)=k_{1}(k_{2}+\sigma_{i}(t))^{\mu}$ with $\mu>1,k_{1},k_{2}\geq 1$ and $c_{i}(t)\geq 1$, $\Sigma_{i}(t)\geq 1$. By selecting the Lyapunov function $\tilde{V}(\vartheta_{t},t)=V(\vartheta_{t},t)$, it can be deduced that $\mathfrak{L}\tilde{V}(\vartheta_{t},t)=\sum_{i=1}^{M}r_{i}[\eta_{1}^{l}\Sigma_{i}(t)\sigma_{i}(t)+2\eta_{2}^{l}c_{i}(t)-2\eta_{2}^{l}\psi_{i}]\dot{c}_{i}(t)+\eta_{1}^{l}\sum_{i=1}^{M}r_{i}c_{i}(t)\Sigma_{i}(t)\mathfrak{L}\sigma_{i}(t)+\sum_{i=M+1}^{N}s_{i-M}[\eta_{1}^{f}\Sigma_{i}(t)\sigma_{i}(t)+2\eta_{2}^{f}c_{i}(t)-2\eta_{2}^{f}\psi_{i}]\dot{c}_{i}(t)+\eta_{1}^{f}\sum_{i=M+1}^{N}s_{i-M}c_{i}(t)\Sigma_{i}(t)\mathfrak{L}\sigma_{i}(t)$. By the similar steps of the above proof of Theorem IV.1, the m.s. and a.s. consensus of (1) can be solved by the protocol (2) with $\Sigma_{i}(t)=k_{1}(k_{2}+\sigma_{i}(t))^{\mu}$, for $\mu>1,k_{1},k_{2}\geq 1$ and $c_{i}(0)\geq 1$, $\Sigma_{i}(0)\geq 1$.
Furthermore, for the case of $\mu=1$ and $\gamma=0$, the control protocol (2) can be written as
$\begin{split}&u_{i}(t)=k_{1}(k_{2}c_{i}(t)+\sigma_{i}(t))\mathcal{K}\xi_{i}(t),\\ &\dot{c}_{i}(t)=\xi_{i}^{T}(t)\Gamma\xi_{i}(t),\end{split}$ (22)
where $k_{1},k_{2}>0$, $\xi_{i}(t)=\sum_{j\in N_{i}}a_{ij}(x_{i}(t)-x_{j}(t))$ and $\sigma_{i}(t)=\xi_{i}^{T}(t)P\xi_{i}^{T}(t)$. Then we have the following theorem.
Theorem IV.2.
Suppose the directed graph $\tilde{\mathcal{G}}$ contains a spanning tree. The m.s. and a.s. consensus of (1) can be solved by the protocol (22) with $\mathcal{\mathcal{K}}=-B^{T}P$, and $\Gamma=PBB^{T}P$, where $P$ is the solution to the SARE (9).
Proof.
Substituting the control protocol (22) into (1) yields the following closed-loop system like (4)
$\begin{split}d\vartheta(t)=&(I_{N}\otimes A+k_{1}(k_{2}\tilde{C}(t)+\tilde{\sigma}(t))\mathcal{L}\otimes B\mathcal{K})\vartheta(t)dt\\ &+(I_{N}\otimes C)\vartheta(t)dw(t).\end{split}$ (23)
Then, according to the graph decomposition in Lemma IV.1, we have
$\begin{split}d\xi_{l}(t)\!=&(I_{M}\!\otimes\!A\!+\!(k_{1}(k_{2}\tilde{C}_{l}(t)\!+\!\tilde{\sigma}_{l}(t))\mathcal{L}_{11})\!\otimes\!B\mathcal{K})\xi_{l}(t)dt\!\!\\ &\!+\!(I_{M}\!\otimes\!C)\xi_{l}(t)dw(t),\end{split}$ (24)
and
$\begin{split}d\xi_{\!f\!}(t)\!=&(I_{N\!-\!M}\!\otimes\!A\!+\!(k_{1}(k_{2}\tilde{C}_{\!f\!}(t)\!+\!\tilde{\sigma}_{\!f\!}(t))\mathcal{L}_{22})\!\otimes\!B\mathcal{K})\xi_{\!f\!}(t)dt\!\!\!\!\!\!\!\\ &+(k_{1}(k_{2}\tilde{C}_{l}(t)\!+\!\tilde{\sigma}_{l}(t))\mathcal{L}_{21}\otimes B\mathcal{K})\xi_{l}(t)dt\\ &+(I_{N\!-\!M}\!\otimes\!C)\xi_{f}(t)dw(t).\end{split}$ (25)
Note that the Lyapunov functional $V(\vartheta_{t},t)$ in Theorem IV.1 satisfies $V(\vartheta_{t},t)\leq\sum_{i=1}^{M}r_{i}[\eta_{1}^{l}k_{1}(2k_{2}c_{i}(t)+\sigma_{i}(t))\sigma_{i}(t)+\eta_{2}^{l}(c_{i}(t)-\psi_{i})^{2}]+\sum_{i=M+1}^{N}s_{i-M}[\eta_{1}^{f}k_{1}(2k_{2}c_{i}(t)+\sigma_{i}(t))\sigma_{i}(t)+\eta_{2}^{f}(c_{i}(t)-\psi_{i})^{2}]$ with $\mu=1$. Then, we choose Lyapunov functional
$\begin{split}V_{1}&(\vartheta_{t},t)\\ =&\sum_{i=1}^{M}\!r_{i}\!\Big[\eta_{1}^{l}k_{1}(2k_{2}c_{i}(t)\!+\!\sigma_{i}(t))\sigma_{i}(t)\!+\!\eta_{2}^{l}(c_{i}(t)\!-\!\psi_{i})^{2}\Big]\\ &\!\!+\!\!\!\!\sum_{i=M+1}^{N}\!\!\!\!\!s_{i-M}\!\Big[\eta_{1}^{f}k_{1}(2k_{2}c_{i}(t)\!+\!\sigma_{i}(t))\sigma_{i}(t)\!+\!\eta_{2}^{f}(c_{i}(t)\!-\!\psi_{i})^{2}\Big],\end{split}$
where $\eta_{1}^{l},\eta_{2}^{l},\eta_{1}^{f},\eta_{2}^{f}>0$, $\psi_{i}$ are positive constants to be determined. According to the functional $\mathrm{It\hat{o}}$ formula, it can be deduced that $dV_{1}(\vartheta_{t},t)$ has the following form
$\begin{split}dV_{1}&(\vartheta_{t},t)\!=\!\mathfrak{L}V_{1}(\vartheta_{t},t)dt+\frac{\partial V_{1}(\vartheta_{t},t)}{\partial\vartheta}(I_{N}\!\otimes\!C)\vartheta(t)dw(t),\end{split}$
where $\mathfrak{L}V_{1}(\vartheta_{t},t)=\sum_{i=1}^{M}r_{i}\big[(\eta_{k_{2}}^{l}\sigma_{i}(t)+2\eta_{2}^{l}c_{i}(t)-2\eta_{2}^{l}\psi_{i})\dot{c}_{i}(t)\big]+\sum_{i=1}^{M}r_{i}\big[\big(\eta_{k_{2}}^{l}c_{i}(t)+\eta_{k_{1}}^{l}\sigma_{i}(t)\big)\mathfrak{L}\sigma_{i}(t)\big]+\sum_{i=M+1}^{N}s_{i-M}\big[(\eta_{k_{2}}^{f}\sigma_{i}(t)+2\eta_{2}^{f}c_{i}(t)-2\eta_{2}^{f}\psi_{i})\dot{c}_{i}(t)\big]+\sum_{i=M+1}^{N}s_{i-M}\big[\big(\eta_{k_{2}}^{f}c_{i}(t)+\eta_{k_{1}}^{f}\sigma_{i}(t)\big)\mathfrak{L}\sigma_{i}(t)\big]$ with $\eta_{k_{1}}^{l}=2\eta_{1}^{l}k_{1}$, $\eta_{k_{2}}^{l}=2\eta_{1}^{l}k_{1}k_{2}$, $\eta_{k_{1}}^{f}=2\eta_{1}^{f}k_{1}$, and $\eta_{k_{2}}^{f}=2\eta_{1}^{f}k_{1}k_{2}$. Denote $\tilde{\psi}=\mathrm{diag}\{\psi_{1},...,\psi_{N}\}$. According to Lemma IV.1, $\tilde{\psi}$ can be partitioned into $\tilde{\psi}_{l}=\mathrm{diag}\{\psi_{1},...,\psi_{M}\}$ and $\tilde{\psi}_{f}=\mathrm{diag}\{\psi_{M+1},...,\psi_{N}\}$. Let $\mathcal{K}=-B^{T}P$ and $\Gamma=PBB^{T}P$. Note that $\sum_{i=1}^{M}\sigma_{i}(t)=\xi_{l}^{T}(t)(I_{M}\otimes P)\xi_{l}(t)$ and $\xi_{l}(t)=(\mathcal{L}_{11}\otimes I_{n})\vartheta_{l}(t)$. By the definition of $\sigma_{i}(t)$, (24), (25) and using Lemma IV.1 and IV.2, we have
$\begin{split}&\sum_{i=1}^{M}\!r_{i}\big[\big(\eta_{k_{2}}^{l}c_{i}(t)\!+\!\eta_{k_{1}}^{l}\sigma_{i}(t)\big)\mathfrak{L}\sigma_{i}(t)\big]\\ &=\xi_{l}^{T}\!(t)[R(\eta_{k_{2}}^{l}\tilde{C}_{l}(t)\!+\!\eta_{k_{1}}^{l}\tilde{\sigma}_{l}(t))\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\xi_{l}(t)\!\!\!\!\\ &\ \ \ -\!2\eta_{1}^{l}\xi_{l}^{T}(t)\big[(k_{1}k_{2}\tilde{C}_{l}(t)+k_{1}\tilde{\sigma}_{l}(t))^{2}\tilde{\mathcal{L}}_{11}\otimes\Gamma\big]\xi_{l}(t),\end{split}$
and
$\begin{split}&\sum_{i=1}^{M}r_{i}(\eta_{k_{2}}^{l}\sigma_{i}(t)+2\eta_{2}^{l}c_{i}(t)-2\eta_{2}^{l}\psi_{i})\dot{c}_{i}(t)\\ &=\xi_{l}^{T}(t)[R(\eta_{k_{2}}^{l}\tilde{\sigma}_{l}(t)\!+\!2\eta_{2}^{l}\tilde{C}_{l}(t)\!-\!2\eta_{2}^{l}\tilde{\psi}_{l})\otimes\Gamma]\xi_{l}(t).\end{split}$
Note that $\sum_{i=M+1}^{N}\sigma_{i}(t)=\xi_{f}^{T}(t)(I_{N-M}\otimes P)\xi_{f}(t)$. Then, using Lemma IV.1, IV.3, and the elementary inequality: $2a^{T}Ob\leq\iota a^{T}Oa+\frac{1}{\iota}b^{T}Ob$, for any positive definite matrix $O\in\mathbb{R}^{n\times n}$, $a,b\in\mathbb{R}^{n}$ and $\iota>0$, we have
$\begin{split}&\sum_{i=M+1}^{N}\!s_{i-M}\big[\big(\eta_{k_{2}}^{f}c_{i}(t)\!+\!\eta_{k_{1}}^{f}\sigma_{i}(t)\big)\mathfrak{L}\sigma_{i}(t)\big]\\ &=\!\xi_{f}^{T}\!(t)[S(\eta_{k_{2}}^{f}\!\tilde{C}_{f}(t)\!+\!\eta_{k_{1}}^{f}\!\tilde{\sigma}_{f}(t))\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\xi_{f}(t)\!\!\!\!\\ &\ \ \ -2\eta_{1}^{f}\xi_{f}^{T}(t)\big[(k_{1}k_{2}\tilde{C}_{f}(t)+k_{1}\tilde{\sigma}_{f}(t))^{2}\tilde{\mathcal{L}}_{22}\otimes\Gamma\big]\xi_{f}(t)\\ &\ \ \ -\!2\eta_{1}^{f}\xi_{f}^{T}\!(t)\big[(k_{1}k_{2}\tilde{C}_{l}(t)\!+\!k_{1}\tilde{\sigma}_{l}(t))\\ &\quad\quad\quad\quad\quad\quad\cdot(k_{1}k_{2}\tilde{C}_{f}(t)\!+\!k_{1}\tilde{\sigma}_{f}(t))S\mathcal{L}_{21}\!\otimes\!\Gamma\big]\xi_{l}(t)\\ &\leq\!\xi_{f}^{T}\!(t)[S(\eta_{k_{2}}^{f}\!\tilde{C}_{f}(t)\!+\!\eta_{k_{1}}^{f}\!\tilde{\sigma}_{f}(t))\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\xi_{f}(t)\!\!\!\!\\ &\ \ \ -2\eta_{1}^{f}\xi_{f}^{T}(t)\big[(k_{1}k_{2}\tilde{C}_{f}(t)+k_{1}\tilde{\sigma}_{f}(t))^{2}\tilde{\mathcal{L}}_{22}\otimes\Gamma\big]\xi_{f}(t)\\ &\ \ \ +2\eta_{1}^{f}\iota\xi_{f}^{T}\!(t)[(k_{1}k_{2}\tilde{C}_{f}(t)\!+\!k_{1}\tilde{\sigma}_{f}(t))^{2}\!\otimes\!\Gamma]\xi_{f}(t)\\ &\ \ \ +2\eta_{1}^{f}\frac{\sigma_{\max}^{2}(S\mathcal{L}_{21})}{\iota}\xi_{l}^{T}\!(t)[(k_{1}k_{2}\tilde{C}_{l}(t)\!+\!k_{1}\tilde{\sigma}_{l}(t))^{2}\!\otimes\!\Gamma]\xi_{l}(t),\end{split}$
and
$\begin{split}&\sum_{i=M+1}^{N}\!s_{i-M}\big[(\eta_{k_{2}}^{f}\sigma_{i}(t)\!+\!2\eta_{2}^{f}c_{i}(t)\!-\!2\eta_{2}^{f}\psi_{i})\dot{c}_{i}(t)\big]\\ &=\xi_{f}^{T}(t)[S(\eta_{k_{2}}^{f}\tilde{\sigma}_{f}(t)\!+\!2\eta_{2}^{f}\tilde{C}_{f}(t)\!-\!2\eta_{2}^{f}\tilde{\psi}_{f})\otimes\Gamma]\xi_{l}(t).\end{split}$
Denote $\Theta^{l}_{t}=\eta_{k_{2}}^{l}\tilde{C}_{l}(t)+\eta_{k_{1}}^{l}\tilde{\sigma}_{l}(t)$ and $\Theta^{f}_{t}=\eta_{k_{2}}^{f}\tilde{C}_{f}(t)+\eta_{k_{1}}^{f}\!\tilde{\sigma}_{f}(t)$. Substituting the above inequalities into $\mathfrak{L}V_{1}(\vartheta_{t},t)$ yields
$\begin{split}\mathfrak{L}&V_{1}(\vartheta_{t},t)\\ \leq&\xi_{l}^{T}\!(t)[R\Theta^{l}_{t}\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\xi_{l}(t)\\ &+\xi_{f}^{T}\!(t)[S\Theta^{f}_{t}\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\xi_{f}(t)\\ &-2\tilde{\eta}_{l}\xi_{l}^{T}\!(t)\big[(k_{1}k_{2}\tilde{C}_{l}(t)\!+\!k_{1}\tilde{\sigma}_{l}(t))^{2}\!\otimes\!\Gamma\big]\xi_{l}(t)\\ &-2\tilde{\eta}_{f}\xi_{f}^{T}(t)\big[(k_{1}k_{2}\tilde{C}_{f}(t)+k_{1}\tilde{\sigma}_{f}(t))^{2}\otimes\Gamma\big]\xi_{f}(t)\\ &+\xi_{l}^{T}(t)[R(\eta_{k_{2}}^{l}\tilde{\sigma}_{l}(t)\!+\!2\eta_{2}^{l}\tilde{C}_{l}(t)\!-\!2\eta_{2}^{l}\tilde{\psi}_{l})\!\otimes\!\Gamma]\xi_{l}(t)\\ &+\xi_{f}^{T}(t)[S(\eta_{k_{2}}^{f}\tilde{\sigma}_{f}(t)\!+\!2\eta_{2}^{f}\tilde{C}_{f}(t)\!-\!2\eta_{2}^{f}\tilde{\psi}_{f})\otimes\Gamma]\xi_{l}(t),\end{split}$
where $\tilde{\eta}_{l}=\eta_{1}^{l}\lambda_{2}(\tilde{\mathcal{L}}_{11})-\eta_{1}^{f}\frac{\sigma_{\max}^{2}(S\mathcal{L}_{21})}{\iota}$ and $\tilde{\eta}_{f}=\eta_{1}^{f}(\lambda_{2}(\tilde{\mathcal{L}}_{22})-\iota)$. Choose parameters $\eta_{1}^{l},\eta_{1}^{f},\iota$ to ensure $\tilde{\eta}_{l}>0$ and $\tilde{\eta}_{f}>0$. Let $\eta_{2}^{l}=\eta_{1}^{l}k_{1}k_{2}^{2}$ and $\eta_{2}^{f}=\eta_{1}^{f}k_{1}k_{2}^{2}$. Using the elementary inequality: $2a^{T}Ob\leq\iota a^{T}Oa+\frac{1}{\iota}b^{T}Ob$, for any positive definite matrix $O\in\mathbb{R}^{n\times n}$, $a,b\in\mathbb{R}^{n}$ and $\iota>0$, ones get
$\begin{split}\mathfrak{L}&V_{1}(\vartheta_{t},t)\\ \leq&\xi_{l}^{T}\!(t)[R\Theta^{l}_{t}\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC\!+\!k_{2}\Gamma)]\xi_{l}(t)\\ &\!\!\!\!+\!\xi_{f}^{T}\!(t)[S\Theta^{f}_{t}\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC\!+\!k_{2}\Gamma)]\xi_{f}(t)\\ &\!\!\!\!-\!\xi_{l}^{T}\!(t)\big[(2\tilde{\eta}_{l}(k_{1}k_{2}\tilde{C}_{l}(t)\!+\!k_{1}\tilde{\sigma}_{l}(t))^{2}\!+\!2\eta_{2}^{l}\tilde{\psi}_{l}R)\!\otimes\!\Gamma\big]\xi_{l}(t)\\ &\!\!\!\!-\!\xi_{f}^{T}(t)\big[(2\tilde{\eta}_{f}(k_{1}k_{2}\tilde{C}_{f}(t)\!+\!k_{1}\tilde{\sigma}_{f}(t))^{2}\!+\!2\eta_{2}^{f}\tilde{\psi}_{f}S)\otimes\Gamma\big]\xi_{f}(t)\\ \leq&\xi_{l}^{T}\!(t)[R\Theta^{l}_{t}\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC\!+\!k_{2}\Gamma)]\xi_{l}(t)\\ &\!\!\!\!+\!\xi_{f}^{T}\!(t)[S\Theta^{f}_{t}\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC\!+\!k_{2}\Gamma)]\xi_{f}(t)\\ &\!\!\!\!-\!2\xi_{l}^{T}\!(t)\Big[\sqrt{\tilde{\eta}_{l}\eta_{2}^{l}\tilde{\psi}_{l}R}\Big(k_{1}k_{2}\tilde{C}_{l}(t)\!+\!k_{1}\tilde{\sigma}_{l}(t)\Big)\!\otimes\!\Gamma\big]\xi_{l}(t)\\ &\!\!\!\!-\!2\xi_{f}^{T}(t)\big[\sqrt{\tilde{\eta}_{f}\eta_{2}^{f}\tilde{\psi}_{f}S}\Big(k_{1}k_{2}\tilde{C}_{f}(t)+k_{1}\tilde{\sigma}_{f}(t)\Big)\otimes\Gamma\Big]\xi_{f}(t).\end{split}$
Let $\psi_{i}\geq\frac{(\eta_{1}^{l})^{2}\max\{r_{i}\}(1+k_{2})^{2}}{\tilde{\eta}_{l}\eta_{2}^{l}}$ for $i=1,\dots,M$ and $\psi_{i}\geq\frac{(\eta_{1}^{f})^{2}\max\{s_{i-M}\}(1+k_{2})^{2}}{\tilde{\eta}_{f}\eta_{2}^{f}}$ for $i=M+1,\dots,N$. Then, we have
$\begin{split}\mathfrak{L}V_{1}(\vartheta_{t},t)\leq&-\!\xi_{l}^{T}\!(t)[R\Theta^{l}_{t}\otimes I_{n}]\xi_{l}(t)\!-\!\xi_{f}^{T}\!(t)[S\Theta^{f}_{t}\otimes I_{n}]\xi_{f}(t)\\ \leq&-\!\varepsilon\xi_{l}^{T}\!(t)\Big[R\Big(\eta_{k_{2}}^{l}\tilde{C}_{l}(t)\!+\!\frac{\eta_{k_{1}}^{l}}{2}\tilde{\sigma}_{l}(t)\Big)\otimes P\Big]\xi_{l}(t)\\ &-\!\varepsilon\xi_{f}^{T}\!(t)\Big[S\Big(\eta_{k_{2}}^{f}\tilde{C}_{f}(t)\!+\!\frac{\eta_{k_{1}}^{f}}{2}\tilde{\sigma}_{f}(t)\Big)\otimes P\Big]\xi_{f}(t)\\ \leq&-\varepsilon V_{1}(\vartheta_{t},t)+\varphi(t),\end{split}$
where $A^{T}P+PA-PBB^{T}P+C^{T}PC+I_{n}=0$, $\varepsilon=\frac{1}{\lambda_{\max}(P)}$, and $\varphi(t)=\eta_{2}^{l}\varepsilon\sum_{i=1}^{M}r_{i}(c_{i}(t)-\psi_{i})^{2}+\eta_{2}^{f}\varepsilon\sum_{i=M+1}^{N}s_{i-M}(c_{i}(t)-\psi_{i})^{2}$. Then, according to Lemma III.1 and the proof of Theorem IV.1, we can prove the existence and uniqueness, as well as m.s. and a.s. stability of the solution to (23). Then, by the definition of $\xi(t)$, the m.s. and a.s. consensus can be solved. ∎
Remark IV.2.
Theorem IV.1 and IV.2 designed fully distributed protocols to solve m.s. and a.s. consensus for stochastic MASs under directed graphs. Different from the existing adaptive protocol designs [24, 26] for stochastic MASs with undirected graphs, Theorem IV.1 and IV.2 considered the case of directed graphs and proposed a fully distributed control protocol that can be applied to solve the m.s. and a.s. consensus under directed graphs.
Furthermore, two time-varying gains $c_{i}(t)$ and $\Sigma_{i}(t)$ are incorporated in the adaptive strategy (2). These two terms serve distinct functions in facilitating stochastic consensus: $c_{i}(t)$ is utilized to estimate global information, while $\Sigma_{i}(t)$ addresses the asymmetrical interactions characteristic of directed graphs. Additionally, owing to the asymmetry inherent in the Laplacian matrix of a directed graph, we have adopted a Lyapunov function that diverges from that used in the case of undirected graphs, and utilized the properties of M-matrix to address this issue.
Remark IV.3.
In fact, the parameter selection criteria based on the Riccati equation remain fundamentally similar to the existing results. Specifically, when the stochastic noise term in the stochastic MASs (1) vanishes, the SARE (9) degenerates into a standard algebraic Riccati equation $A^{T}P+PA-PBB^{T}P+\mathcal{Q}=0$ ($\mathcal{Q}=I_{n}$) applicable to deterministic MASs. This special case aligns perfectly with the results in [35], demonstrating the theoretical consistency between our approach and previous works.
Remark IV.4.
Theorems IV.1 and IV.2 established a unified and flexible framework for the design of fully distributed protocol in stochastic environments under directed topology. The core of the design lies in the auxiliary time-varying gain function $\Sigma_{i}(t)$, whose structure (parameterized by $k_{1}$, $k_{2}$, $\mu$, and the time-varying term $c_{i}(t)$, $\sigma_{i}(t)$) introduces significant generality in protocol design.
This generality enables our adaptive protocol to encompass several notable consensus designs for deterministic MASs under directed topologies as special cases, demonstrating broader applicability. Specifically:
  • The feedback structure proposed in [35] emerges as a special case of Remark IV.1 with $\mu=3$ and $k_{1}=k_{2}=1$.
  • The adaptive gain design presented in [14] can be obtained from Theorem IV.2 by setting $k_{1}=k_{2}=1$.
Therefore, the protocols developed in [35] and [14] can be viewed as particular parameterizations within our proposed framework (2). This generalization not only unifies existing approaches but also provides enhanced flexibility for performance tuning and robustness improvement in the presence of complex stochastic dynamics.
IV-B Fully distributed consensus control for stochastic MASs with the undirected topology
The parameter $\Sigma_{i}(t)$ in the protocol (2) is introduced to account for asymmetric interactions in the network; however, since undirected graphs exhibit symmetric adjacency relations by definition, $\Sigma_{i}(t)$ becomes superfluous and can be safely omitted in such case. Assume that the undirected graph $\bar{\mathcal{G}}$ is connected. Therefore, the eigenvalues of $\mathcal{L}$ for the undirected graph $\bar{\mathcal{G}}$ are real and nonnegative, denoted by $\lambda_{1}(\mathcal{L})=0$ and $0<\lambda_{2}(\mathcal{L})\leq\ldots\leq\lambda_{N}(\mathcal{L})$.
Then, we have the following theorem for the case of undirected topology.
Theorem IV.3.
Suppose the undirected graph $\bar{\mathcal{G}}$ is connected. The m.s. and a.s. consensus of (1) can be solved by the protocol (2) with $\Sigma_{i}(t)=1$, $\gamma=0$, $\mathcal{K}=-B^{T}P$, and $\Gamma=PBB^{T}P$, where $P$ is the solution to the SARE (9).
Proof.
The proof is in Appendix A1. ∎
Although the directed graph protocol discussed with $\gamma=0$ in the above theorem is theoretically applicable to undirected graphs, we exploit the advantageous structural properties of undirected graphs to develop an enhanced distributed control protocol. In the following, we consider a fully distributed control protocol with tunable parameters $\gamma$ specifically tailored for undirected graphs, which demonstrates superior convergence performance compared to the previously established protocol for the case of directed graphs.
The control protocol (2) can be written as
$\begin{split}&u_{i}(t)=c_{i}(t)\mathcal{K}\xi_{i}(t),\\ &\dot{c}_{i}(t)=e^{\gamma t}\xi_{i}^{T}(t)\Gamma\xi_{i}(t),\end{split}$ (26)
where $\xi_{i}(t)=\sum_{j\in N_{i}}a_{ij}(x_{i}(t)-x_{j}(t))$, $c_{i}(t)$ is an adaptive gain functinon, $\mathcal{K}\in\mathbb{R}^{m\times n}$ and $\Gamma\in\mathbb{R}^{n\times n}$ are the feedback gain matrices, and $\gamma$ is a non-negative constant to be determined. Under the assumption that the undirected graph $\bar{\mathcal{G}}$ is connected, we can show that the m.s. and a.s. consensus can be solved by (26).
Theorem IV.4.
Suppose the undirected graph $\bar{\mathcal{G}}$ is connected. The m.s. and a.s. consensus of (1) can be solved by the protocol (26) with $\mathcal{K}=-B^{T}P$ and $\Gamma=PBB^{T}P$, where $P$ is the solution to the SARE (9) and $\frac{1}{\lambda_{\max}(P)}\leq\gamma<\frac{3}{2\lambda_{\max}(P)}$.
Proof.
Substituting the control protocol (26) into (1) yields the following closed-loop system
$\begin{split}dx(t)=&(I_{N}\otimes A+\tilde{C}(t)\mathcal{L}\otimes B\mathcal{K})x(t)dt\\ &+(I_{N}\otimes C)x(t)dw(t),\end{split}$ (27)
where $\tilde{C}(t)=\mathrm{diag}\{c_{1}(t),...,c_{N}(t)\}$. Let $\vartheta_{i}(t)=x_{i}(t)-\frac{1}{N}\sum_{j=1}^{N}x_{j}(t)$. Denote $\vartheta(t)=[(I_{N}-\frac{1}{N}\mathbf{1}_{N}\mathbf{1}_{N}^{T})\otimes I_{n}]x(t)=[\vartheta_{1}^{T}(t),...,\vartheta_{N}^{T}(t)]^{T}$. Then, the closed-loop system (27) can be written as follows
$\begin{split}d\vartheta(t)=&(I_{N}\otimes A+\tilde{C}(t)\mathcal{L}\otimes B\mathcal{K})\vartheta(t)dt\\ &+(I_{N}\otimes C)\vartheta(t)dw(t).\end{split}$ (28)
Firstly, we choose the Lyapunov function
$\begin{split}V_{3}(\vartheta_{t},t)=\vartheta^{T}(\mathcal{L}\otimes P)\vartheta+\sum_{i=1}^{N}e^{-\gamma t}(c_{i}(t)-\psi_{i})^{2},\end{split}$ (29)
where $P>0$ and $\psi_{i}>0$. Applying the functional $\mathrm{It\hat{o}}$ formula to (29), we have
$\begin{split}dV_{3}&(\vartheta_{t},t)=\mathfrak{L}V_{3}(\vartheta_{t},t)dt+2\vartheta^{T}(\mathcal{L}\otimes PC)\vartheta(t)dw(t),\end{split}$
where $\mathfrak{L}V_{3}(\vartheta_{t},t)$ is defined as
$\begin{split}\mathfrak{L}V_{3}&(\vartheta_{t},t)=\vartheta^{T}(t)[\mathcal{L}\otimes(A^{T}P+PA+C^{T}PC)]\vartheta(t)\\ &+2\vartheta^{T}(t)[\mathcal{L}\tilde{C}(t)\mathcal{L}\otimes PB\mathcal{K}]\vartheta(t)\\ &+2\sum_{i=1}^{N}e^{-\gamma t}(c_{i}(t)-\psi_{i})\dot{c}_{i}(t)-\gamma\sum_{i=1}^{N}e^{-\gamma t}(c_{i}(t)-\psi_{i})^{2}.\end{split}$
Let $\mathcal{K}=-B^{T}P$ and $\Gamma=PBB^{T}P$. Then, we can obtain
$\begin{split}\mathfrak{L}V_{3}&(\vartheta_{t},t)=\vartheta^{T}(t)[\mathcal{L}\otimes(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\vartheta(t)\\ &-2\vartheta^{T}(t)[\mathcal{L}\tilde{C}(t)\mathcal{L}\otimes PBB^{T}P]\vartheta(t)\\ &+2\sum_{i=1}^{N}(c_{i}(t)-\psi_{i})\xi_{i}^{T}(t)PBB^{T}P\xi_{i}(t)\\ &-\gamma\sum_{i=1}^{N}e^{-\gamma t}(c_{i}(t)-\psi_{i})^{2}.\end{split}$
Denote $\varepsilon=\frac{1}{\lambda_{\max}(P)}$. Let $\psi_{i}>\frac{1}{2\lambda_{2}(\mathcal{L})}$ and $\gamma\geq\varepsilon$. Note that $A^{T}P+PA-PBB^{T}P+C^{T}PC+I_{n}=0$. Thus, Substituting (37) into $\mathfrak{L}V_{3}(\vartheta_{t},t)$ yields
$\begin{split}\mathfrak{L}&V_{3}(\vartheta_{t},t)\\ \leq&\vartheta^{T}\!(t)[\mathcal{L}\otimes(A^{T}P+PA-2\tilde{\psi}\lambda_{2}(\mathcal{L})PBB^{T}P)]\vartheta(t)\\ &+\!\vartheta^{T}\!(t)[\mathcal{L}\!\otimes\!(C^{T}\!PC)]\vartheta(t)\!-\!\gamma\sum_{i=1}^{N}\!e^{-\gamma t}(c_{i}(t)\!-\!\psi_{i})^{2}\\ \leq&-\vartheta^{T}(t)[\mathcal{L}\otimes I_{n}]\vartheta(t)-\gamma\sum_{i=1}^{N}e^{-\gamma t}(c_{i}(t)-\psi_{i})^{2}\\ \leq&-\varepsilon V_{3}(\vartheta_{t},t).\end{split}$ (30)
Next, we will prove the existence and uniqueness, m.s. stability, and a.s. stability of the solution to (28) in the following three steps.
Step 1: Similar to Step 1 in the proof of Theorem IV.1, it can be seen that the drift and diffusion terms of closed-loop stochastic system (28) satisfy the local Lipschitz condition. According to Lemma III.1, the existence and uniqueness of the solution to (28) can be proved.
Step 2: By the functional $\mathrm{It\hat{o}}$ formula for $V_{3}(\vartheta_{t},t)$, we have
$\begin{split}&\mathrm{E}(e^{\delta(\tau_{h}\wedge t)}V_{3}(\vartheta_{\tau_{h}\wedge t},\tau_{h}\wedge t))-V_{3}(\vartheta_{0},0)\\ &=\mathrm{E}\int_{0}^{\tau_{h}\wedge t}e^{\delta s}[\delta V_{3}(\vartheta_{s},s)+\mathfrak{L}V_{3}(\vartheta_{s},s)]ds,\end{split}$
where $\delta>0$. Note that $\check{V}_{3}(\vartheta)\leq V_{3}(\vartheta_{t},t)$, where $\check{V}_{3}(\vartheta)=\vartheta^{T}(\mathcal{L}\otimes P)\vartheta$. Substituting the inequality $\check{V}_{3}(\vartheta)\leq V_{3}(\vartheta_{t},t)$ and (30) into the above equation yields
$\begin{split}&\mathrm{E}(e^{\delta(\tau_{h}\wedge t)}\check{V}_{3}(\vartheta_{\tau_{h}\wedge t})-C_{3}\\ &\leq(\delta-\varepsilon)\mathrm{E}\int_{0}^{\tau_{h}\wedge t}e^{\delta s}V_{3}(\vartheta_{s},s)ds,\end{split}$
where $C_{3}=\check{V}_{3}(\vartheta_{0})=\vartheta^{T}(0)(\mathcal{L}\otimes P)\vartheta(0)$. Let $\delta=\varepsilon$, then $\mathrm{E}(e^{\delta(\tau_{h}\wedge t)}\check{V}_{3}(\vartheta_{\tau_{h}\wedge t}))\leq C_{3}$. Thus, let $h\to\infty$, then we can obtain by using the Fatou lemma that
$\begin{split}\mathrm{E}(e^{\delta t}\check{V}_{3}(\vartheta_{t}))\leq C_{3}.\end{split}$ (31)
That is, $\mathrm{E}(\check{V}_{3}(\vartheta_{t}))\leq C_{3}e^{-\delta t}$, which implies
$\begin{split}\lim_{t\to\infty}\sup\frac{1}{t}\log\mathrm{E}(\check{V}_{3}(\vartheta_{t}))\leq-\delta.\end{split}$ (32)
This together with the definition of $\check{V}_{3}$ also produces
$\lim_{t\to\infty}\sup\frac{1}{t}\log(\mathrm{E}|\vartheta(t)|^{2})\leq-\delta,$ (33)
That is, the m.s. exponential stability of the closed-loop system (28) follows. According to the definition of $\vartheta(t)$, the m.s. consensus can be solved.
Step 3: According to the functional $\mathrm{It\hat{o}}$ formula, it can be seen that
$\begin{split}&e^{\delta t}V_{3}(\vartheta_{t},t)-V_{3}(\vartheta_{0},0)\\ &=\int_{0}^{t}e^{\delta s}[\delta V_{3}(\vartheta(s),s)+\mathfrak{L}V_{3}(\vartheta_{s},s)]ds+M_{v3}(t),\end{split}$
where $M_{v3}(t)$ is a local martingale with $M_{v3}(0)=0$. Then, similar to (31), it can be deduced that
$\begin{split}e^{\delta t}\check{V}_{3}(\vartheta_{t})\leq C_{3}+M_{v3}(t),\end{split}$
Using the non-negative semi-martingale convergence theorem [38], ones have $\lim_{t\to\infty}\sup[e^{\delta t}\check{V}_{3}(\vartheta_{t})]<\infty$ a.s. Hence,
$\begin{split}\lim_{t\to\infty}\sup\frac{1}{t}\log(\check{V}_{3}(\vartheta_{t}))\leq-\delta,a.s.\end{split}$ (34)
Then, (34) together with the definition of $\check{V}_{3}$ yields
$\lim_{t\to\infty}\sup\frac{1}{t}\log(|\vartheta(t)|)\leq-\frac{\delta}{2}\ a.s.$
That is, the a.s. exponential stability of the closed-loop system (28) follows. According to the definition of $\vartheta(t)$, the a.s. consensus can be solved.
Furthermore, the boundedness of input can be further elucidated below.
$\begin{split}|u(t)|=&|(\int_{0}^{t}e^{\gamma s}\vartheta^{T}(s)R_{1}\vartheta(s)ds+C_{1})R_{2}\vartheta(t)|\\ \leq&\lambda_{\max}(R_{1}R_{2})\int_{0}^{t}e^{\gamma s}|\vartheta(s)|^{2}ds|\vartheta(t)|+C_{1}R_{2}|\vartheta(t)|,\end{split}$
where $R_{1}=\mathcal{L}^{2}\otimes PBB^{T}P$, $C_{1}=\vartheta^{T}(0)R_{1}\vartheta(0)$ and $R_{2}=\mathcal{L}\otimes B^{T}P$. Note that $\lim_{t\to\infty}\sup\frac{1}{t}\log(|\vartheta(t)|)\leq-\frac{\delta}{2}\ a.s.$ Hence, for any $0<\kappa<\frac{\delta}{2}$, one can find a positive random variable $\varpi=\varpi(\kappa)$ such that $|\vartheta(t)|\leq\varpi e^{-(\frac{\delta}{2}-\kappa)t}$ for all $t\geq 0$. Thus, it can be obtained that
$\begin{split}|u(t)|\leq&\lambda_{\max}(R_{1}R_{2})\varpi^{3}\int_{0}^{t}e^{(\gamma-\delta+2\kappa)s}ds\ e^{-(\frac{\delta}{2}-\kappa)t}\\ &+C_{1}R_{2}\varpi e^{-(\frac{\delta}{2}-\kappa)t}\\ \leq&\frac{\lambda_{\max}(R_{1}R_{2})\varpi^{3}}{\gamma-\delta+2\kappa}(e^{(\gamma-\frac{3\delta}{2}+3\kappa)t}-e^{-(\frac{\delta}{2}-\kappa)t})\\ &+C_{1}R_{2}\varpi e^{-(\frac{\delta}{2}-\kappa)t}.\end{split}$
Note that $\delta=\varepsilon=\frac{1}{\lambda_{\max}(P)}$ and $\gamma<\frac{3}{2\lambda_{\max}(P)}$. For sufficiently small $\kappa$, $\gamma-\frac{3\delta}{2}+3\kappa<0$. Therefore, it can be deduced that the input $u(t)$ is bounded. ∎
Moreover, according to Theorem IV.3 in the preceding analysis, the adjustable parameter $\gamma$ is only required to satisfy $0<\gamma<\frac{3}{2\lambda_{\max}(P)}$ to guarantee the boundedness of input. The additional constraint $\gamma\geq\frac{1}{\lambda_{\max}(P)}$ in Theorem IV.4 is introduced solely to derive the explicit exponential convergence rate of the stochastic MASs.
Remark IV.5.
The inclusion of the exponential term $e^{\gamma t}$ introduces only a scalar operation per time step, whose computational cost is negligible relative to the distributed quadratic-form updates already inherent to the algorithm. Compared with the protocol presented in Theorem IV.3, the proposed scheme incorporates an adjustable parameter $\gamma$, thereby enhancing the flexibility of the control design. Moreover, it can be shown that an improved convergence rate is attainable through appropriate tuning of this parameter.
Remark IV.6.
Theorem IV.4 proposed a fully distributed protocol for stochastic MASs to solve m.s. and a.s. consensus under undirected graphs. In comparison to the existing fully distributed stochastic consensus results[24, 26], we derive specific exponential estimates for the convergence rates of m.s. and a.s. consensus based on the proof of Theorem IV.4, that is, the relationship between the exponential convergence rate and system parameters is established.
Remark IV.7.
The consensus algorithm for undirected graphs proposed in this work is inspired by the protocol presented in [25]. While both address the consensus problem under similar stochastic settings, our contributions extend beyond the existing works, establishing a more comprehensive theoretical foundation:
1) Theoretical completeness: In contrast to [25], which primarily focuses on algorithm design and convergence analysis, we provide a rigorous proof of existence and uniqueness of solutions for the closed-loop stochastic differential system. This is accomplished by establishing verifiable sufficient conditions (Lemma III.1) within a stochastic Lyapunov framework.
2) Practical feasibility: We establish the uniform boundedness of the proposed protocol under the parameter constraint $\gamma<\frac{3}{2\lambda_{\max}(P)}$. This result guarantees that the control inputs remain bounded throughout the evolution of the system, a critical property for practical implementation that was not explicitly analyzed in [25].
3) Topological generality: The core theoretical contribution lies in the development of a fully distributed consensus framework applicable to directed graphs (Theorems IV.1 and IV.2). By incorporating the auxiliary time-varying gain function $\Sigma_{i}(t)$, our design eliminates the reliance on global network information and the assumption of topological symmetry, which is typically required for undirected graph cases like [25]. This extension significantly broadens the applicability to realistic networked systems with asymmetric information exchange.
Consequently, while inspired by [25], this work not only consolidates the theoretical basis for undirected topologies but also provides a novel and practical fully distributed approach for the more general directed case in stochastic environments.
Remark IV.8.
The protocol discussed in this section is a node-based adaptive protocol. Compared to edge-based adaptive protocols, this approach generally exhibits lower dimensionality since the number of edges in a connected graph typically exceeds that of its nodes. However, it is worth noting that edge-based adaptive protocols possess their own advantages, such as being more applicable to switching topology because the selection of Lyapunov function does not rely on the Laplacian matrix of a fixed topology. This aspect represents a potential avenue for future research as indicated in [24].
Remark IV.9.
The framework established in this paper can be naturally extended to nonlinear systems. While the current analysis focuses on linear dynamics with path-dependent adaptive feedback, the core analytical approach, including the construction of Lyapunov functions, the application of functional $\mathrm{It\hat{o}}$ formula, and the use of existence and uniqueness, essentially does not rely on the linear dynamics. For nonlinear dynamics satisfying appropriate growth conditions (such as, local Lipschitz condition and polynomial growth), a similar fully distributed protocol design and consensus analysis can be carried out for this class of stochastic nonlinear MASs. Such extensions will be discussed in future work.
V Simulations
In this section, two simulation examples are given to verify the effectiveness of the proposed theoretical results.
Consider the stochastic MASs (1) composed of six agents, where $A=\begin{bmatrix}-0.5&0.1\\ 0&-20\end{bmatrix}$, $B=\begin{bmatrix}0\\ 1\end{bmatrix}$, and $C=\begin{bmatrix}0&0\\ 0&6.5\end{bmatrix}$. According to the SARE (9), we can obtain $P=\begin{bmatrix}1&0.0047\\ 0.0047&0.9046\end{bmatrix}$. Then, it can be deduced that $\mathcal{K}=\begin{bmatrix}-0.0047&-0.9046\end{bmatrix}$, and $\Gamma=\begin{bmatrix}0&0.0042\\ 0.0042&0.8182\end{bmatrix}$.
V-A The case of directed topology
The interaction topology is modeled as a directed graph $\tilde{\mathcal{G}}$, as shown in Figure 2.
Figure 1: The directed graph
Figure 2: The undirected graph
The control input $u_{i}(t)$ adopts the adaptive control protocol (22). The initial states for each agent are randomly chosen from [2,2][-2,2]. We choose $k_{1}=k_{2}=1$ and $c_{i}(0)=1$, i=1,,6i=1,...,6. The revolutions of adaptive gains $c_{i}(t)$, i=1,,6i=1,...,6 are shown in Figure 4, implying that $c_{i}(t)$ will converge to a finite positive constant. The relative states $x_{i1}(t)-x_{11}(t)$, $x_{i2}(t)-x_{12}(t)$, i=1,,6i=1,...,6 of one sample path and the behaviors of the m.s. relative states $E\|x_{i1}(t)-x_{11}(t)\|^{2}$}, $E\|x_{i2}(t)-x_{12}(t)\|^{2}$}, i=1,,6i=1,...,6 of $10^{2}$ sample paths are shown in Figure 4, indicating that a.s. and m.s. consensus can be achieved.
Figure 3: The adaptive gains
Figure 4: The relative states errors and m.s. relative state errors
V-B The case of undirected topology
The interaction topology is modeled as an connected undirected graph $\bar{\mathcal{G}}$, as shown in Figure 2. The control input $u_{i}(t)$ adopts the adaptive control protocol (26). The initial states for each agent are randomly chosen from [2,2][-2,2]. By Theorem IV.4, we can choose $\gamma=1$. Let $c_{i}(0)=1$, i=1,,6i=1,...,6. The revolutions of adaptive gains $c_{i}(t)$ are shown in Figure 6. Considering the relative state of one sample path for each agent $x_{i1}(t)-x_{11}(t)$ and $x_{i2}(t)-x_{12}(t)$, we have Figure 6, which indicates that states of the six agents tend to be a.s. consensus over time. For m.s. consensus analysis, we generate $10^{2}$ sample paths. Then, considering the behaviors of the m.s. relative states $E\|x_{i1}(t)-x_{11}(t)\|^{2}$} and $E\|x_{i2}(t)-x_{12}(t)\|^{2}$}, we obtain Figure 6, which demonstrates that the six agents reach m.s. consensus.
Figure 5: The adaptive gains
Figure 6: The relative states errors and m.s. relative state errors
[…]
Figure 7: The m.s. relative state errors for the case of $\gamma=0,0.5,1$
VI Conclusion
The design of a unified fully distributed consensus protocol for a class of stochastic MASs under directed and undirected graphs are investigated. The existence and uniqueness of solutions for path-dependent and highly nonlinear closed-loop stochastic MASs under fully distributed protocols are first explored. For the case of directed graphs, a unified fully distributed control protocol is designed for the first time to solve m.s. and a.s. consensus for stochastic MASs under directed graphs. Then, for the case of undirected graphs, we develop an enhanced fully distributed protocol and derive explicit exponential convergence rates for both m.s. and a.s. consensus, precisely characterizing their dependence on system parameters. Future research directions include extending these results to MASs with compound disturbances and investigating more challenging scenarios involving heterogeneous nonlinear dynamics and time-varying communication topologies.
A1: The proof of Theorem IV.3
Proof.
Let $\Sigma_{i}(t)=1$. As mentioned in Theorem IV.1, the closed-loop MASs can be written as follows
$\begin{split}d\vartheta(t)=&(I_{N}\otimes A+\tilde{C}(t)\mathcal{L}\otimes B\mathcal{K})\vartheta(t)dt\\ &+(I_{N}\otimes C)\vartheta(t)dw(t).\end{split}$ (35)
Then, we choose Lyapunov function
$\begin{split}V_{2}(\vartheta_{t},t)=\vartheta^{T}(\mathcal{L}\otimes P)\vartheta+\frac{1}{2}\sum_{i=1}^{N}(c_{i}(t)-\psi_{i})^{2},\end{split}$ (36)
where $P>0$ and $\psi_{i}>0$. Applying the functional $\mathrm{It\hat{o}}$ formula to (36), we have
$\begin{split}dV_{2}&(\vartheta_{t},t)\!=\!\mathfrak{L}V_{2}(\vartheta_{t},t)dt+2\vartheta^{T}\!(t)(\mathcal{L}\otimes PC)\vartheta(t)dw(t),\end{split}$
where $\mathfrak{L}V_{2}(\vartheta_{t},t)$ is defined as
$\begin{split}\mathfrak{L}V_{2}&(\vartheta_{t},t)=\vartheta^{T}(t)[\mathcal{L}\otimes(A^{T}P+PA+C^{T}PC)]\vartheta(t)\\ &+2\vartheta^{T}(t)[\mathcal{L}\tilde{C}(t)\mathcal{L}\otimes PB\mathcal{K}]\vartheta(t)+(c_{i}(t)-\psi_{i})\dot{c}_{i}(t).\end{split}$
Let $\mathcal{K}=-B^{T}P$ and $\Gamma=PBB^{T}P$. Then, we can obtain
$\begin{split}\mathfrak{L}V_{2}&(\vartheta_{t},t)=\vartheta^{T}(t)[\mathcal{L}\otimes(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\vartheta(t)\\ &-2\vartheta^{T}(t)[\mathcal{L}\tilde{C}(t)\mathcal{L}\otimes PBB^{T}P]\vartheta(t)\\ &+\sum_{i=1}^{N}(c_{i}(t)-\psi_{i})\xi_{i}^{T}(t)PBB^{T}P\xi_{i}(t).\end{split}$
Note that $\xi_{i}(t)=\sum_{j\in N_{i}}(x_{i}(t)-x_{j}(t))=\sum_{j\in N_{i}}(\vartheta_{i}(t)-\vartheta_{j}(t))$. Then, we have
$\begin{split}&\sum_{i=1}^{N}(c_{i}(t)-\psi_{i})\xi_{i}^{T}(t)PBB^{T}P\xi_{i}(t)\\ &=\vartheta^{T}(t)[\mathcal{L}(\tilde{C}(t)-\tilde{\psi})\mathcal{L}\otimes PBB^{T}P]\vartheta(t),\end{split}$ (37)
where $\tilde{\psi}=\mathrm{diag}\{\psi_{1},...,\psi_{N}\}$. Denote $\varepsilon=\frac{1}{\lambda_{\max}(P)}$. Let $\psi_{i}>\frac{1}{\lambda_{2}(\mathcal{L})}$. Note that $A^{T}P+PA-PBB^{T}P+C^{T}PC+I_{n}=0$. Thus, Substituting (37) into $\mathfrak{L}V_{2}(\vartheta_{t},t)$ yields
$\begin{split}\mathfrak{L}&V_{2}(\vartheta_{t},t)\\ \leq&\vartheta^{T}(t)[\mathcal{L}\otimes(A^{T}P+PA-\tilde{\psi}\lambda_{2}(\mathcal{L})PBB^{T}P)]\vartheta(t)\\ &+\vartheta^{T}(t)[\mathcal{L}\otimes(C^{T}PC)]\vartheta(t)\\ \leq&-\vartheta^{T}(t)[\mathcal{L}\otimes I_{n}]\vartheta(t)\\ \leq&-\varepsilon V_{2}(\vartheta_{t},t)+\bar{\varphi}(t),\end{split}$ (38)
where $\bar{\varphi}(t)=\frac{\varepsilon}{2}\sum_{i=1}^{N}(c_{i}(t)-\psi_{i})^{2}$. By Lemma III.1 and the similar steps of Theorem IV.1, the m.s. and a.s. consensus of (1) can be solved by the protocol (2) with $\Sigma_{i}(t)=1$ for the case of undirected topology. ∎